### v1

# Understanding In-Context Learning with a Pelican Soup Framework

Abstract
Many existing theoretical analyses of in-context learning for natural language processing are based on latent variable models that leaves gaps between theory and practice. We aim to close these gaps by proposing a theoretical framework, the Pelican Soup Framework. In this framework, we introduce (1) the notion of a common sense knowledge base, (2) a general formalism for natural language classification tasks, and the notion of (3) meaning association. Under this framework, we can establish a $\mathcal{O}(1/T)$ loss bound for in-context learning, where $T$ is the number of example-label pairs in the demonstration. Compared with previous works, our bound reflects the effect of the choice of verbalizers and the effect of instruction tuning. An additional notion of atom concepts makes our framework possible to explain the generalization to tasks unseen in the language model training data. Finally, we propose a toy setup, Calcutec, and a digit addition task that mimics types of distribution shifts a model needs to overcome to perform in-context learning. We also experiment with GPT2-Large on real-world NLP tasks. Our empirical results demonstrate the efficacy of our framework to explain in-context learning.
In-Context Learning, Pelican Soup Framework
1 Introduction
Large language models (LM) have demonstrated the ability to perform downstream natural language processing (NLP) classification tasks via in-context learning (ICL) (Chowdhery et al., 2022). That is, they can predict the label of an input by conditioning on the demonstration, i.e., a sequence of input-label pairs (Brown et al., 2020), even when the verbalizers (labels present in the demonstration) are semantically irrelevant to the task, e.g., foo/bar instead of negative/positive (Wei et al., 2023). To understand why LMs trained with nonstructural general text can perform ICL, a line of works base their theoretical analysis on the assumption that the data is from a latent variable model (Xie et al., 2022; Zhang et al., 2023; Wang et al., 2023b).
However, the assumptions on a latent variable model in these works leave a gap between theory and practice. Xie et al. (2022) assumes that the general text for training LMs is from a hidden Markov model, which is an oversimplification of natural language. Zhang et al. (2023) propose a more general latent variable model for the data generation process. Although their model is more realistic than a hidden Markov model, their analyses are based on the assumption that there is a data generation process that generates both the general text for LM training and the demonstration for in-context learning. As demonstrations have a format very different from general text, a question to be answered is whether such a generation process really exists. Additionally, considering the diversity of the semantics natural language can express, it is likely that the training data does not cover all possible values in the latent variable’s space. It requires more characterizations on the latent space to discuss the generalization to semantics unseen in the training data.
Besides, previous theoretical analyses do not fully answer some questions either (Xie et al., 2022; Zhang et al., 2023; Hahn & Goyal, 2023). For example, what is the effect of the choice of verbalizers on ICL performance? How does the distribution of downstream task examples affect ICL? How instruction-tuning helps ICL (Chung et al., 2022)?
To answer these questions, we propose a theoretical framework, the Pelican Soup Framework. We introduce the notion of a commonsense knowledge base (KB) and define the data generation process as the process that generates documents consistent with the KB. We also introduce a formalism for NLP classification tasks. This notion and formalism instantiate a data generation process that satisfies the assumption made by Zhang et al. (2023). With slight adaption on the Zhang et al. (2023), we establish a similar bound for ICL loss under our framework. Additionally, because we see the latent variable as the association between the verbalizers and their meaning, our bound can reflect the effect of the choice of verbalizers. It also reflects how instruction tuning affects ICL performance.
If we further assume that all latent meaning can be represented with logic formula over a finite set of atom concepts as suggested by cognitive science theories Fodor (1975, 2008); Piantadosi (2021), we can characterize the generalization to the unseen meaning during inference time. Having this assumption instantiates a setup compatible with the theory for the emergence of complex skills by (Arora & Goyal, 2023), and we can thus apply their analyses on generalization to our framework. We also discuss our results with the results by Hahn & Goyal (2023), where they bound ICL loss with the description length of the ground truth input-label mapping function.
Inevitably, whether our analyses are meaningful depends on how deep models such as transformers (Vaswani et al., 2017) generalize. Specifically, as most of the previous theoretical analyses do, we show that a perfect LM that estimates the next token distribution perfectly according to the training data distribution can do ICL. However, when doing ICL, LMs are conditioned on the demonstration, which has a low probability mass in the training data. LMs can do ICL only if it can generalize well under this distribution shift from training data to ICL demonstrations.
As the generalization of LMs is difficult to analyze theoretically, we conduct experiments to show empirical evidence for our framework. We construct a synthetic dataset, Calcutec (§4), and a synthetic task, the addition task (§5), to mimic the distribution shifts LMs need to overcome. With Calcutec, we show that LMs can do ICL regardless of the distributional shifts and observe the efficacy of chain-of-thoughts. With the addition task, we observe that larger models generalize better under a certain type of distributional shift, which is aligned with the observation about model scaling in the real world Wei et al. (2022). Finally, in §6, we experiment with natural language and show that using pronouns as verbalizers can lead to comparable performance (with GPT-2 Large), which is aligned with our framework where LMs learn ICL ability by modeling meaning association of words.
2 The Pelican Soup Framework
2.1 Motivation: Pelican Soup
The Pelican Soup game inspires our framework. It is a game involving a puzzle master who has a story in their mind. The game participants aim to recover the story by asking the puzzle master yes/no questions. An observation is that once the participants recover the story, they can answer any questions about the story. Therefore, the story has a similar role as a latent variable defining the input-output mapping, and the yes/no questions are similar to the demonstrations for in-context learning. We include an example in Appendix A.
Given the above observation, we can study in-context learning by considering why humans can solve Pelican Soup riddles. We conjecture that this is because the person who makes the story and the ones who solve the riddle share the same (or similar) commonsense (McCarthy, 1960) about logical relationships among things in this world (Schank & Abelson, 1988). This inspires us to introduce the notion of a commonsense knowledge base in our framework.
2.2 Training Data Distribution
Our theory framework is based on the interrelations between the semantics of sentences. As suggested by the long history of linguistics studies that aim at representing sentence semantics with formal languages (Carnap et al., 1968; Bresnan & Bresnan, 1982; Steedman, 1987, 1996; Sag et al., 1999), such as first-order logic (Frege et al., 1879; Peirce, 1883), it is reasonable to make a general assumption on language:
Assumption 2.1 (Sentence).
We assume that a sentence in a language is a sequence of words such that humans can determine whether one sentence entails another sentence or whether two sentences contradict each other.
Motivated by the Pelican Soup game and the early formulation of AI (McCarthy, 1960; Croft, 1993), we introduce the notion of a commonsense knowledge base and use it to restrict the possible combination of sentences that can co-occur in a paragraph:
Assumption 2.2 (Knowledge base).
There is a knowledge base $\mathrm{KB}$, which is a set of sentences, such that any paragraph with non-zero probability mass is a set of sentences such that for any two sentences $x_{1},x_{2}$ in the paragraph, $x_{1}\wedge\mathrm{KB}$ does not contradict with $x_{2}$.
It is natural to assume the existence of this knowledge base. This knowledge base may include statements describing interrelations among the concepts represented by words, as cognitive psychologists have suggested word semantics are based on these interrelations (Siskind, 1996; Murphy, 2004). For example, the knowledge base may contain a rule: “If X moved, then the position of X changed” because the semantic of the verb “move” is embodied in its implication on “position”. Based on this knowledge base, humans can do step-by-step reasoning with language. Humans also avoid writing paragraphs that include contradictory sentences.
We also observe that language allows us to arbitrarily associate concepts to an entity or a pronoun, which we refer to as reference. For example, when the pronoun “she” or the human name “Emily” is present in a paragraph, it is associated with a person that can be described with a set of sentences. The usage of the reference is dependent on the meaning it is associated with and is consistent within its context. For example, if “she” is associated with the sentence “a person who has a house”, then the sentence “she has no property” will have 0 probability mass. Moreover, when we want to refer to “the person who has a house” instead of repeating the sentence again, we use “she” for an abbreviation (when there is no ambiguity).
We can see abstract adjectives such as “good” and “bad” as references too, as their meaning also depends on the context. However, their meaning may not be as variable as pronouns. We reflect this difference with a prior distribution for the meaning a reference is associated with.
For simplicity, we only consider single-word references and assume that such association is consistent throughout a document. We write down an assumption formally:
Assumption 2.3 (Reference).
A set of words $\Gamma\subset\Sigma$ serve as references in the language. For every document in the training data, some $r\in\Gamma$ can be associated with a meaning represented as a set of sentences $Z_{r}$, which follows a prior distribution $\Pr(Z_{r})$. Any $z\in Z_{r}$, $z$ present in the document can be replaced by $r$.
Finally, we assume a document is composed of a set of paragraphs where some references are present:
Assumption 2.4 (Document).
A document is a concatenation of paragraphs separated with a delimiter $d$ (e.g. a blank line). Also, there is a set of reference $S\subset\Gamma$ such that in each of the paragraphs, a $r\in S$ is present.
2.3 A Formalism for NLP Classification Tasks
With the commonsense knowledge base $\mathrm{KB}$ defined in Assumption 2.2, we propose a simple formalism: For any objective or prescriptive NLP classification task (Rottger et al., 2022) that classifies an input $x$ to one of $|\mathcal{Y}|$ classes, it can be described with some instructions $u$ and some descriptions $v_{i}$ for $i=1,2,\cdots,|\mathcal{Y}|$ such that $x$ is labeled as $y_{i}$ if and only if
$\mathrm{KB}\wedge t\wedge x\models v_{i}.$ (1)
That is, $x$ belongs to class $y_{i}$ if and only if based on the commonsense rules in $\mathrm{KB}$ and the instruction $t$, $x$ entails $v_{i}$. Note that because we can rewrite Eq. 1 as $\mathrm{KB}\wedge x\models u\to z_{i}$, we can combine the task instructions with the description of each class $v_{i}$ for $i=1,2\cdots,|\mathcal{Y}|$. Therefore, we can represent any NLP classification tasks as $\langle v_{1},v_{2},\cdots,v_{|\mathcal{Y}|}\rangle$.
[…]
Assumption 2.5 (Downstream Task).
NLP classification tasks can be defined with the formulation described above. Namely, given any classification task with classes $\mathcal{Y}$, we can define it with the descriptions $\{v_{y}\}_{y\in\mathcal{Y}}$. We further assume that the descriptions of different classes contradict with each other, so every input belongs to only one class.
2.4 Bounding ICL Loss
Our framework is quite general; by adapting and combining the analyses by Zhang et al. (2023) and Hahn & Goyal (2023), we have the following theorem:
Theorem 2.6 (Average ICL Loss).
Let the description of a classification task be $\{z_{y}\}_{y\in\mathcal{Y}}$ and $z^{*}$ be the association between the task description $\{v_{y}\}_{y\in\mathcal{Y}}$ and the verbalizers $\Gamma_{\mathcal{Y}}$ used for ICL. Let $K$ be the constraints used for decoding, $\dot{g}$ be the event where a document belongs to a certain genera. Let $S_{t}=x_{1},r_{2},d,x_{2},r_{2},d\cdots,x_{t},r_{t},d$, where $r_{t}$ is the verbalizer that is associated with the label of $x_{t}$ and $d$ is the delimiter. We have for any positive integer $T$, the average cross-entropy loss of ICL is bounded as:
$\begin{split}&-\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1},K)\\ &\leq\begin{aligned} &-\frac{1}{T}\log\Pr(z^{*},\dot{g}|K)\\ &-\frac{1}{T}\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},\dot{g},S_{t-1},K)\\ &-\frac{1}{T}\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K)}{\Pr(x_{t}|S_{t-1},K)}\end{aligned}\end{split}$ (2)
When the last two terms on the right-hand side are non-negative, Eq. 2 shows the average cross-entropy loss of ICL converges to 00 in $\mathcal{O}(1/T)$. We discuss the three terms in the right-hand side below.
The second term becomes 00 if we set $K$ as the constraint that the next two tokens of $S_{t-1},x_{t}$ must be a verbalizer and the delimiter for all $t$. This is because Assumption 2.2 ensures that $x_{t}$ does not conflict with $r_{t}$ and Assumption 2.5 ensures that $x_{t}$ conflict with verbalizers other than $r_{t}$.
We then look at the last term. This term is 00 when $x_{t}$ is conditionally independent to $z^{*}$ as assumed by Zhang et al. (2023). However, this may be an over-simplification because, in natural language, the transition from $x_{t}$ to its next token depends on the content of $x_{t}$. This assumption may actually be unnecessary for convergence because $x_{t}$ is an example from a downstream task related to $z^{*}$; it is likely that
$\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K)\geq\Pr(x_{t}|S_{t-1},K),$
which implies that this term is non-negative, and we can thus ignore this term.
More rigorously, we can show the following corollary:
Corollary 2.7 (Expected Average ICL Loss).
Let $\dot{g}$ represent a set of documents whose paragraphs are conditionally independent to each other given $z^{*}$, i.e. $\Pr(x_{1},d,x_{2},d,\cdots,x_{T},d|z^{*},\dot{g})=\prod_{t=1}^{T}\Pr(x_{t},d|z^{*},\dot{g})$. If the downstream task data distribution $\mathcal{D}_{X}$ follow $\Pr(x|z^{*},\dot{g})$, then we can bound the average ICL cross-entropy loss over the downstream task as:
$\begin{split}&\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\left[-\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1},K)\right]\\ &\leq-\frac{1}{T}\log\Pr(z^{*},\dot{g}|K).\end{split}$ (3)
The right-hand side of Eq. 3 characterizes the convergence rate and reflects the difficulty of doing ICL. If $z^{*},\dot{g},K$ are independent, then we can see this term is proportional to $\Pr(z^{*})$. This implies that when the association between label description and the verbalizer is uncommon in the training data (e.g., associating “positive” to “This movie is bad.”), doing ICL is more difficult.
Eq. 2 also allows us to analyze the scenario where we do not constrain the next token of $S_{t-1},x_{t}$ to be a verbalizer while decoding. We can replace $\dot{g}$ with $\ddot{g}$ that represents the documents satisfying $\dot{g}$ and the constraint $K$ (i.e., verbalizers always follow $x$). [^1] Although $\ddot{g}$ may seem unnatural, this genere of documents corresponds to the PCFG structure assumed in Hahn & Goyal (2023). This ensures the second term of Eq. 2 to be 00. The cost of using $\ddot{g}$ makes the first term in bound bigger because $\Pr(\ddot{g}|z^{*})\leq\Pr(\dot{g}|z^{*})$. This reflects that doing ICL without constraining the next token is harder.
$\Pr(\ddot{g}|z^{*})$ may also be related to instruction tuning. The training examples for instruction tuning are input-output pairs following some format, such as having an instruction (e.g., “Label the example as positive if …”) at the beginning and a prompt after each input (e.g. “[example]. The sentiment of this comment is”). Because for examples that follow the special format, the next token after the prompt is always a verbalizer, these examples belong to the genera $\ddot{g}$. Having these examples in the training set would increase $\Pr(\ddot{g}|z^{*})$ and thus make the ICL loss bound converge faster. This explains why instruction tuning helps ICL.
2.5 Extending the Framework with Formal Language
Assuming a latent model poses a dilemma: Language can encode various meanings, so assuming that the latent space is finite is not reasonable, or the space must be very large. However, if the latent space is infinite or is very large, it is possible that the limited training data does not cover the whole space. Without any assumption on the latent space (e.g., the relation between the states in the space), it is impossible to discuss the generalization to unseen latent states. Thus, we provide an extension to our framework.
We can address this dilemma by extending our framework with an assumption stronger than Assumption 2.1:
Assumption 2.8 (Meaning representation).
There exists (1) a finite set of atom concepts $\Omega$ and (2) a function $f$ that can map any sentence in language to its meaning represented with a logical formula with operands in $\Omega$ such that for any two sentences $s_{1},s_{2}$, the logical relation between $s_{1}$ and $s_{2}$ is the same as $f(s_{1})$ and $f(s_{2})$.
This assumption is aligned with the cognitive psychology studies that hypothesize the existence of a set of mental tokens (Fodor, 1975, 2008) and a recent study (Piantadosi, 2021) showing that semantics can be encoded with the combination of only a few symbols.
This assumption suggests that if we have the parsing function $f$, solving NLP tasks only requires a finite-length program that can do logical reasoning by manipulating logical symbols according to logical induction rules. If a deep model can learn this program, then it can perform a task even if this task is not in the training data. This assumption of a finite $\Omega$ also instantiates the concept of language skills by Arora & Goyal (2023), and their theoretical results are thus applicable.
2.6 Relating to Function Description Length
When there is no decoding constrains, $\Pr(r_{t},d|x_{t},z^{*},S_{t-1})$ is related to the difficulty of the example. To see this, we need an additional assumption:
Assumption 2.9.
In some documents in the training data, the paragraphs are constituted with steps in a logical induction process with some steps randomly dropped.
This kind of document may be pervasive in the training data. Essays arguing some claims are one example. To be convincing, these essays should proceed like a proving process that induces their conclusions. Documents describing a series of events can be another example, as events follow commonsense and develop progressively.
With this assumption and some regularity assumptions on the data distribution, we can have
$\Pr(r_{t},d|x_{t},z^{*},S_{t-1})\leq c\cdot\ell(x_{t}),$ (4)
where $\ell(x_{t})$ is the number of reasoning steps required to solve the task, and $c$ is a constant This $\ell(x_{t})$ corresponds to the description length of the function that maps the inputs to their label in the loss bound by Hahn & Goyal (2023) (more discussion in Appendix D).
Train
x57 x56 x64 r3 $\to$ x79 , r1 x57 $\to$ x58 , x90 x58 $\to$ r3 , r3 r1 $\to$ x20 , … ; x80 x66 x63 x83 x1 $\to$ x82 , x80 x82 $\to$ r1 , … , x64 x80 $\to$ x54 .
ICL
x44 x67 x34 x62 $\to$ r2 ; x55 x38 x50 x48 $\to$ r1 ; x21 x59 x57 x86 $\to$ r2 ; x55 x76 x84 x99 $\to$
CoT
x44 x67 x34 x62 $\to$ x16 , x34 x62 $\to$ x99 , x99 x16 $\to$ r1 ; x77 x34 x62 x97 $\to$ x12 … ; x21 x59 x57 x86 $\to$ x69 , x59 x57 $\to$ x75 , x69 x75 $\to$ r2 ; x55 x76 x84 x99 $\to$
Table 1: Calcutec examples for training, in-context learning (ICL), and chain-of-thought (CoT).
3 Distribution Shifts from Language Modeling to In-context Learning
In addition to the theoretical analyses, an important aspect of understanding ICL is to understand how models generalize. The bounds we have established in § 2.4 are based on the analyses of the marginal distribution conditioned on the input $S_{t-1},x_{t}$. Because the distribution of prefix $S_{t-1},x_{t}$ is different from the distribution of general text, a model needs to overcome several types of distribution shifts to predict the marginal distribution precisely. We characterize a few of them below.
Format mismatch.
The demonstration is in the format $x_{t},r_{t},d$, where $x_{t}$ follows the downstream task distribution. However, in the LM training data, sequences in this format are unlikely to happen frequently.
Verbalizer Mismatch.
The verbalizers appear very frequently in $S_{t-1}$. However, the verbalizer may not have as high frequency in the training data distribution.
Unseen Reference-Meaning Association.
When performing ICL, the verbalizers are associated with descriptions of the classes for the task. However, the training data may not include this kind of association.
We empirically investigate these types of distribution shifts in the next section.
4 Experimenting with Synthetic Datasets
As a concrete instance of our framework, we present a toy setup, Calcutec, which includes a data generation process for LM training data and a process for downstream tasks. Calcutec follows our framework and mimics the distribution shifts LMs need to overcome as discussed in §3. We will use this setup to inspect models’ generalization ability.
4.1 Calcutec
Setup.
Following the assumption in §2, we construct a pseudo-language with the following components:
  • Logic model: We use a subset of propositional logic as our logic model. We only consider Horn clauses (Horn, 1951), i.e., formulas in the form $A\wedge B\to C$. This allows us to efficiently generate proof with time complexity linear to the size of the knowledge base (Dowling & Gallier, 1984).
  • Atom concepts: We use 100 variables as our set of atom concepts $\Sigma$.
  • $\mathrm{KB}$: For each variable $\sigma\in\Sigma$, there are 5 formulas in the knowledge base whose consequence (the literals after $\to$ in a horn clause) is $\sigma$, while the antecedents (the literals before $\to$) are sampled from $\Sigma\backslash\{\sigma\}$ uniformly at random.
  • We have a set $\Gamma=\{r_{i}\}_{i=1}^{4}$ representing the reference words described in Assumption 2.3.
Training Dataset.
Following Assumption 2.4, a document is a concatenation of paragraphs separated by delimiter “;” and ends with “.”. In our synthetic language model training dataset, each document contains 16 paragraphs. Before starting the generation process, we randomly draw a reference symbol $r_{a}\in\Gamma$ and associate it with two formulas $s_{1},s_{2}$. For simplicity, we make these two formulas single-variable, i.e., $s_{1},s_{2}\in\Sigma$.
Because paragraphs in the real world are sentences ordered in a certain way, we follow Assumption 2.9 and randomly generate a proof as a paragraph. Each paragraph represents the induction process of $P\models g$ for some randomly selected $P\subset\Sigma$ and $g\in\Sigma$ that allows us to use the reference symbol $r$ in the paragraph. Each sentence in the paragraph is a sentence representing a reasoning step. Because we use Horn logic as our logic model, a step is a rule in $\mathrm{KB}$ in the form $a_{1}a_{2}\to b$. We separate the clauses in the sequence with commas. To simulate the fact that documents in the real world always skip some reasoning steps, we further apply some perturbations over the generated paragraphs that drop some reasoning steps with a skip rate $p_{skip}$ (details in Appendix E).  [^2] Models can acquire in-context learning ability even with $p_{skip}=0$ (Figure 6 in thr appendix).
After we generate a document, we replace $s_{1}$ and $s_{2}$ with the reference symbol $r_{a}$. We randomly draw another reference symbol $r_{b}\in\Gamma$ and replace the two most frequent symbols in the document with $r_{b}$. (We provide the pseudo-code Alg. 1 in the appendix.)
Downstream Tasks.
Following the formulation in Assumption 2.5, we define a binary classification task by defining the description $v_{+}$ and $v_{-}$ for the positive and negative classes, respectively. $v_{+}$ and $v_{-}$ are the disjunctions of atom concepts, i.e., in the form of $a_{1}\vee a_{2}\vee\cdots$. We create five downstream tasks using different disjunctions. Each input is a subset of variables in $\Sigma$ from which we ensure that only one of the classes can be induced.
Demonstration.
We represent an input-label pair as $x_{1}x_{2}\cdots\to r$, where $x_{1}x_{2}\cdots$ is the input part and $r\in\{r_{+},r_{-}\}\subset\Gamma$ is a reference symbol serving as a verbalizer.
Chain-of-thought.
A chain-of-thought is in the format same format as the training data, but ends with a reference symbol $r\in\{r_{+},r_{-}\}$, e.g., $x_{1}x_{2}\cdots\to x_{3};x_{3}\cdots x_{4}\to r_{+}$. This chain-of-thought reflects the step-by-step induction process from the inputs to the label.
4.2 Inspecting the Distribution Shifts
Our Calcutec dataset allows us to inspect some types of distribution shifts in § 3.
Format Mismatch.
The training data contains reasoning steps but these steps are absent in the demonstration.
Verbalizer Mismatch.
When we are picking the reference symbols in $\Gamma$, we assign probability mass 45%, 45%, 5%, 5% to $r_{1},r_{2},r_{3},r_{4}$. In this way, we can inspect whether using less frequent verbalizers will lead to worse performance.
Unseen Tasks.
To investigate whether the model can generalize to a new combination of formulas unseen in the training data when we generate our training data, we ensure that the references are never associated with some combinations of atom concepts ($s_{1},s_{2}\in\Sigma$). We then test the trained model on tasks where $v_{+}$ and $v_{-}$ are the disjunctions of unseen combinations. Additionally, while in the training data, each reference symbol is associated with only two atom concepts; we test whether the model can perform in-context learning on tasks where $v_{+}$ and $v_{-}$ are the disjunctions of three atom concepts.
4.3 Experiment Details
We use $p_{skip}=0.25$ in our experiment. We generate 60,000 documents with 16 paragraphs following the process described in § 4. Among them, we use 10k for validation. We train a 6-layer Transformer (Vaswani et al., 2017) model until the loss on the validation set converges. We include additional setups in §G
4.4 Experiment Results and Discussion
(a) Verbalizers = $r_{1},r_{2}$
(b) Verbalizers = $r_{3},r_{4}$
Figure 1: In-context learning accuracy with Calcutec when using different verbalizers ($r_{1},r_{2}$ or $r_{3},r_{4}$). The dotted lines represent the performance of unseen combinations described in §4.2. The colors represent the number of formulas in $\Sigma$ each class ($v_{+}$ or $v_{-}$) is associated with. The main lines represent the average accuracy of 5 tasks. We plot the performance of each task in lighter colors.
$r_{1},r_{2}$ $r_{3},r_{4}$
Task ICL CoT ICL CoT
Single 57.1 91.7 55.6 92.0
Double 53.5 76.3 51.1 77.1
Triple 53.0 73.0 51.7 73.4
Table 2: The 4-shot accuracy of in-context learning (ICL) versus chain-of-thought (CoT) using different verbalizers.
task direct pronoun
SST-2 63.0 65.3
CR 61.7 62.9
MR 59.2 56.7
Subj 51.0 62.2
Table 3: The accuracy of using task-specific templates/verbalizers (direct) (Min et al., 2022a) v.s. using task-agnostic templates/pronouns for 16-shot in-context learning with GPT2-Large.
Training
4 9 2 8 4 6 + 0 8 0 3 5 0 = 0 0 0 0 0 0 ; 0 9 2 8 4 6 + 0 8 0 3 5 0 = 4 0 0 0 0 0 ; 0 0 2 8 4 6 + 0 0 0 3 5 0 = 4 7 1 0 0 0 ; 0 0 0 8 4 6 + 0 0 0 3 5 0 = 4 7 3 0 0 0 ; $\cdots$ 0 0 0 0 0 0 + 0 0 0 0 0 0 = 4 7 3 1 0 7 ;
Testing
8 7 4 0 1 6 + 0 9 2 1 5 0 = 0 0 0 0 0 0 ; 0 0 0 0 0 0 + 0 0 0 0 0 0 =
Table 4: Training and testing examples for the digit addition task.
Figure 1 shows that the model trained with Calcutec can perform in-context learning, indicating that the model can generalize under the distribution shifts described in §3. This evidence supports our Pelican Soup hypothesis. We further inspect the in-context learning performance under the distribution shifts described in §4.2:
  • Infrequent verbalizer: We observe similar performance of using frequent reference symbols ($r_{1},r_{2}$) or infrequent reference symbols ($r_{3},r_{4}$) as the verbalizers.
  • Unseen tasks: Figure 1 shows that the model has similar performance over tasks defined with unseen and unseen combinations of atom concepts (dot lines and solid lines). The models can also generalize to tasks defined with three formulas (green lines).
In sum, the results show that the model can generalize well under several distributional shifts.
We experiment with 4-shot learning using chain-of-thought. Table 2 shows that the model also benefits from chain-of-thought. We conjecture that it is because chain-of-thought has a format more similar to the format of training.
5 Digit Addition Task
In § 4, we show that transformer models can generalize to downstream tasks where the intermediate reasoning steps are missing. As this is related to models’ ability to predict $P(r|x,K)$2.6) when we force the model to skip the intermediate steps, we are curious whether transformer models can still generalize even when more steps are missing. Since in the Calcutec environment, it is difficult to construct a task that requires a long reasoning process without exponentially increasing the size of the atom concept set $\Sigma$, We propose another toy task, the digit addition task, which involves longer reasoning processes.
5.1 Setup
The digit addition task requires the model to perform the addition operation over two $n$-digit numbers. It requires $n$ steps to solve if we solve it digit-by-digit. We generate a training set consisting of digit-by-digit reasoning steps for 100,000 pairs of numbers. Each intermediate step may be dropped at a constant probability $p_{drop}$. After training a Transformer model with the language modeling objective, we test whether the model can generate the final answer without generating the intermediate steps. We experiment with different models of different sizes. From this, we can verify whether models of different sizes can acquire the intuition from the language modeling task.
We show examples of our data in Table 4. Each digit sequence represents a number from the lower digit to the higher digit. The reasoning process in the training set gradually updates both sides of “=” from the lowest digit to the highest digit. As for the testing example, we skip the intermediate steps, making the model complete right-hand side of “=” in the last step. The model needs to acquire intuition from the training set in order to solve the testing samples. (We include a rigorous description in Appendix I.)
Figure 2: The exact accuracy (y-axis, solid points) and digit-level accuracy (y-axis, hollow points) versus validation loss (x-axis) for the 5-digit addition task for five random seeds and different dropping rates $p_{drop}$ (In the appendix, we include the complete results in Figure 11 and the results of the 6-digit addition task in Figure 12.)
5.2 Results and Discussion
We report the exact match accuracy and the digit-level accuracy of models trained with different $p_{drop}$ in Figure 2 with 5 random seeds. A higher accuracy implies the better intuition the model acquires from the sequence modeling task. The results show that three of the four models can achieve perfect accuracy when $p_{drop}$ is as small as $0.3$. It suggests that it is possible that the model can predict the final answer even without intermediate steps, even when examples in the training data are mostly step-by-step. We also observe that larger models tend to have higher and more stable accuracy. When the number of digits is 6 (Figure 12 in the appendix), only the largest model can achieve perfect accuracy. This observation is aligned with the emergence of large language models’ ability.
6 Real-world Evidence
We inspect whether LMs can do ICL with pronouns well because pronouns are reference words frequently associated with different meaning and our framework suggests that LMs learn ICL ability by modeling the association between reference words and their meaning. We thus experiment with the template [input]”, [verbalizer] thought. and use “he”, “she” as the verbalizers. We follow the setup in Min et al. (2022a) and compare the accuracy of the binary classification tasks using GPT2-Large.
Table 3 shows that this task-agnostic template with pronouns is competitive with those task-specific templates. This contradicts with the belief that only larger models can do in-context learning with task-irrelevant verbalizer Wei et al. (2023). It suggests that modeling reference-meaning may indeed contribute LMs’ ICL ability.
7 Related Work
Since Brown et al. (2020) discovered large language models’ in-context learning ability, some theoretical works have attempted to explain how language models acquire this ability. Based on a hidden Markov model (HMM) assumption on the language generation process, Xie et al. (2022) suggested that in-context learning is an implicit Bayesian inference process. Hahn & Goyal (2023) defined the generation process with Compositional Attribute Grammar, which is weaker than the HMM assumption, explaining the in-context learning ability with the minimum description length. Zhang et al. (2023) assumed a more general latent variable model. Arora & Goyal (2023) analyze the emergence of skills based on the scaling law (Hoffmann et al., 2022). While their analysis assumes a set of skills as the atomic elements of NLP tasks, our framework is based on a set of atom concepts.
There were also many empirical studies on the in-context learning ability. Some works focused on the effect of the instruction (Webson & Pavlick, 2022; Lampinen et al., 2022; Jang et al., 2023), while some focused on the examples in the demonstration (Liu et al., 2022; Lu et al., 2022; Sorensen et al., 2022; Min et al., 2022b; Yoo et al., 2022; Ye et al., 2023; Chang & Jia, 2023; Ye et al., 2023; Wang et al., 2023b; Kossen et al., 2023). Shin et al. (2022) found that not all training corpora led to in-context learning ability. Prystawski & Goodman (2023) used synthetic data to suggest that the pretraining dataset’s locality structure contributes to the reasoning steps’ effectiveness. Wang et al. (2023a) studied the reasoning steps in chain-of-thought. Akyürek et al. (2024) formulated ICL as learning a formal language from demonstrations and benchmarked model families.
Some previous work studied in-context learning as a meta-learning-like problem (Chen et al., 2022). Some works focused on the relationships between in-context learning and optimization algorithms (Garg et al., 2022; von Oswald et al., 2022; Akyürek et al., 2023; Fu et al., 2023; Guo et al., 2023). Some works inspected the mechanism of ICL in transformer models (Hendel et al., 2023; Bietti et al., 2023; Todd et al., 2023; Shen et al., 2023; Bai et al., 2023). Chan et al. (2022) studied the properties of dataset distribution that could contribute to the in-context learning ability. Li et al. (2023) provided generalization bounds based on the stability of Transformer models and the distance of downstream tasks. Compared to these works, we focus on how the pretraining data in natural language contributes to the in-context learning ability.
8 Conclusion
In this work, we propose a framework that instantiates a latent variable model for analyzing ICL. Compared with previous works (Xie et al., 2022; Zhang et al., 2023), our latent model better reflects the complexity of language. Under our framework, we establish an ICL loss bond that provides insights into the effect of the choice of verbalizers, downstream task distribution, and instruction tuning. We further provide extensions to the framework that allow us to analyze the generalization to unseen tasks using the results by Arora & Goyal (2023) and relate our bound to the function description length discussed by Hahn & Goyal (2023).
We then present a toy setup, Calcutec, as a concrete instance of our framework. Calcutec mimics the distribution shifts LMs need to overcome to perform ICL. Our empirical results show that transformers trained with autoregressive loss can generalize under these distributional shifts. We additionally use an addition task to show that larger models generalize better, potentially explaining why scaling up model size helps LM performance. Finally, we experiment with real-world NLP tasks and show that GPT-Large can do ICL with pronouns, providing evidence supporting our framework. We expect that our toy datasets can facilitate future research on ICL.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgements
We appreciate Joshua Robinson, Ollie Liu, and Xinyan Velocity Yu for providing valuable feedback on this paper.
Appendix A A Pelican Soup Example
Puzzle master: A men walks into a restaurant and orders pelican soup. After taking a sip, he loses his mind. Why?
Participants: Is it because the soup is not cooked well?
Puzzle master: No.
Participants: Is it because the soup toxic?
Puzzle master: No.
Participants: Does the soup remind him something?
Puzzle master: Yes.
Participants: Did someone cook pelican soup for him?
Puzzle master: Yes.
Participants: Is that person still alive?
Puzzle master: No.
For the sake of aesthetics, we do not include the latent story here. If you are interested, please check it online.
Appendix B Proof of Theorem 2
Let $S_{t}=x_{1},r_{2},x_{2},r_{2},\cdots,x_{t},r_{t}$.
$\Pr(z,g|S_{t},K)=$ $\frac{\Pr(S_{t}|z,g,K)\Pr(z|K)}{\sum_{z}\Pr(S_{t}|z,g,K)\Pr(z,g|K)}$
$=$ $\frac{\Pr(z^{\prime},g|K)\prod_{i=1}^{t}\Pr(x_{t},r_{t},d|z,g,S_{t-1},K)}{\sum_{z^{\prime},g}\Pr(z,g|K)\prod{i=1}^{t}\Pr(x_{t},r_{t},d|z^{\prime},g,S_{t-1},K)}$
$P(x_{t+1},r_{t+1},d|S_{t},K)=$ $\sum_{z,g}\Pr(x_{t+1},r_{t+1},d|z,S_{t},K)\Pr(z,g|S_{t},K)$
$=$ $\frac{\sum_{z,g}\Pr(z,g|K)\prod_{i=1}^{t+1}\Pr(x_{t},r_{t},d|z,g,S_{t-1},K)}{\sum_{z^{\prime}}\Pr(z,g|K)\prod_{i=1}^{t}\Pr(x_{t},r_{t},d|z^{\prime},g,S_{t-1},K)}.$
Thus, it holds that
$-\sum_{t=0}^{T}\log\Pr(x_{t+1},r_{t+1},d|S_{t},K)=$ $-\sum_{t=0}^{T}\begin{aligned} \Big{(}&\log\sum_{z,g}\Pr(z,g|K)\prod_{i=1}^{t+1}\Pr(x_{t},r_{t},d|z,g,S_{t-1},K)\\ &\quad-\log\sum_{z,g}\Pr(z,g|K)\prod_{i=1}^{t}\Pr(x_{t},r_{t},d|z,g,S_{t-1},K)\Big{)}\end{aligned}$
$=$ $-\log\sum_{z,g}\Pr(z,g|K)\prod_{i=1}^{T+1}\Pr(x_{t},r_{t},d|z,g,S_{t-1},K)$
$\leq$ $-\log\Pr(z^{*},\dot{g}|K)\prod_{i=1}^{T+1}\Pr(x_{t},r_{t},d|z^{*},\dot{g},S_{t-1},K)$
$=$ $-\log\Pr(z^{*},\dot{g}|K)-\sum_{i=1}^{T+1}\log\Pr(x_{t},r_{t},d|z^{*},\dot{g},S_{t-1},K)$
$=$ $-\log\Pr(z^{*},\dot{g}|K)-\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},\dot{g},S_{t-1},K)-\sum_{i=1}^{T}\log\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K).$
Thus,
$-\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1},K)$
$\leq$ $-\frac{1}{T}\left(\log\Pr(z^{*},\dot{g}|K)+\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},\dot{g},S_{t-1},K)+\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K)}{\Pr(x_{t}|S_{t-1},K)}\right)$
$+\frac{1}{T}\sum_{i=1}^{T}\log\Pr(d|r_{t},x_{t},S_{t-1},K).$
$\leq$ $-\frac{1}{T}\left(\log\Pr(z^{*},\dot{g}|K)+\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},\dot{g},S_{t-1},K)+\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K)}{\Pr(x_{t}|S_{t-1},K)}\right)$
Appendix C Proof of Corollary 3
The second term in the right-hand side of Eq. 2 is zero when the decoding constrain $K$ is imposed. Therefore, it suffices to prove the last term is non-negative in expectation.
$\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K)}{\Pr(x_{t}|S_{t-1},K)}=$ $\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},K)}{\Pr(x_{t}|S_{t-1},K)}$
$=$ $\sum_{x_{1},x_{2},\cdots,x_{T}}\Pr(x_{t}|z^{*},\dot{g},K)\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},K)}{\Pr(x_{t}|S_{t-1},K)}$
$=$ $\mathrm{KLD}(\Pr(x_{t}|z^{*},\dot{g},K)||\Pr(x_{t}|S_{t-1},K))\geq 0$
Appendix D The Connection between $P(r_{t}|x_{t},z^{*},\ddot{g})$ and Function Description Length by Hahn & Goyal (2023)
Firstly, we make some regularity assumptions: Given a step-by-step reasoning process $\pi=s_{1},s_{2},\cdots,s_{n}$ for the induction process of $P\models Q$, in the training data,
  1. 1.
    each step may be dropped independently to each other with probability $p_{drop}$.
  2. 2.
    $\Pr(s_{i}|P,s_{1},s_{2},\cdots,s_{i-1})>p_{min}$ for all $i\in[n]$.
We first show how we derive Eq. 4: Based on Assumption 2.9,
$\Pr(r_{t}|x_{t},z^{*},\ddot{g})=\sum_{\pi\in\Pi}\Pr(\pi,r_{t}|x_{t},z^{*},\ddot{g})\Pr(\text{$\pi$ is dropped}),$
where $\Pi$ is a set of token sequences representing reasoning steps that induce $r_{t}$ from $x_{t}$. Let $\pi^{*}$ be the shortest proof in $\Pi$, we have
$\log\Pr(r_{t}|x_{t},z^{*},\ddot{g})=$ $\log\sum_{\pi\in\Pi}\Pr(\pi,r_{t}|x_{t},z^{*},\ddot{g})\Pr(\text{$\pi$ is dropped})$
$\geq$ $\log\Pr(\pi^{*},r_{t}|x_{t},z^{*},\ddot{g})\Pr(\text{$\pi^{*}$ isdropped})$
$\geq$ $p_{min}\log\ell(\pi^{*})+p_{drop}\log\ell(\pi^{*}).$
Then we can discuss the connection between $\Pr(r_{t}|x_{t},z^{*},\ddot{g})$ and the function description length by Hahn & Goyal (2023). We can see the dropped reasoning steps in $\pi^{*}$ as the hidden (tree) structure that maps $x_{t}$ to $r_{t}$ as the derivation tree $\tau_{\phi}$ in the bound of Hahn & Goyal (2023). The length of the reasoning steps thus corresponds to the description length of the derivation tree $\mathrm{D}(\tau_{\phi})$.
A major difference between the bound by Hahn & Goyal (2023) and our bound is that their bound has $\mathrm{D}(\tau_{\phi})$ constant to $T$ while our bound has $\sum_{t}\log\Pr(r_{t}|x_{t},z^{*},\ddot{g})$, which potentially grows proportionally to $T$. The cause of this difference is that, Hahn & Goyal (2023) assumes a structure that repetitively applies a function mapping in a document, and the number of repetition is independent to the complexity of the function mapping. In comparison, our framework does not make this assumption.
Appendix E Detailed Gengeration Process of the LM Training Data in Calcutec
We generate a paragraph based on Assumption 2.3 in the following step:
  1. 1.
    We pick a symbol $s$ from the symbols associated with $r_{a}$ uniformly at random.
  2. 2.
    We randomly generate a proof for $\mathrm{KB},P\models g$, where $P\subset\Sigma$ is the premise and $g\in\Sigma$ is the goal of the proof. We ensure that this proof contains the topic $s$.
  3. 3.
    We convert the proof tree to a sequence of proving steps by traversing the proving tree in a topological order with ties broken randomly. Each node in the proof tree corresponds to a rule in $\mathrm{KB}$, so the resulting sequence of proving steps consists of horn clauses in the form $a_{1}a_{2}\to b$. We separate the clauses in the sequence with commas.
  4. 4.
    We rewrite the first step of the proving process to contain the premises of the proof. Specifically, we replace the antecedent in the first formula with the premise $P$. We find that this step is necessary to prevent the language model trained on it from hallucinating irrelevant variables randomly. It is important for our experiment for chain-of-thought, but is not necessary for language models to learn the in-context learning ability.
  Sample $r_{a},r_{b}$ from $\{r_{1},r_{2},r_{3},r_{4}\}$ with probability 0.45, 0.45, 0.05, 0.05.
  Sample topic $S=\{s_{1},s_{2}\}\subset\Sigma$.
  Initialize a document $D$ with empty string.
  for $p=1,2,\dots,n_{par}$ do
     while True do
        Sample $s\in S$.
        Sample a set $X\subset\Sigma$ such that $\bigwedge_{x\in X}x\models s$.
        Run the resolution algorithm to get the set $M=\{m|X\models m\}$.
        Find an extra premise $x^{\prime}$ that can increase the depth of deepest proof tree for $X\models m$.
        Run the resolution algorithm to get the set $M^{\prime}=\{m|X\cup\{x^{\prime}\}\models m\}$.
        if $|M^{\prime}|>\frac{|\Sigma|}{2}$ then
           Reject the sampled $X\cup\{x^{\prime}\}$. {We don’t want a premise that entails everything.}
           Restart the while loop.
        end if
        Sample a $g\in M^{\prime}$ such that the proof tree for $X^{\prime}\models g$ contains $s$ and its depth $>d_{min}$. {We use $d_{min}=4$ in our experiments.}
        Do topological sort to flatten the proof tree and convert it into a string.
        Append the string to $D$.
     end while
  end for
  for $s\in S$ do
     $D\leftarrow$ $D$.replace($s$, $r_{a}$)
  end for
  Let $S^{\prime}=\{s^{\prime}_{1},s^{\prime}_{2}\}\in\Sigma$ be the top-2 frequent non $r_{a}$ symbols in $D$.
  for $s^{\prime}\in S^{\prime}$ do
     $D\leftarrow$ $D$.replace($s^{\prime}$, $r_{b}$)
  end for
Algorithm 1 Pseudo code for the generation process of an Calcutec document used for training.
Appendix F Perturbations in Calcutec
We apply two types of perturbations over the reasoning steps in Calcutec described in §4:
  1. 1.
    Random merge: At probability $p_{merge}$, for every two consecutive clauses where the consequence of the first one is in the antecedents of the second one, say $a_{1}a_{2}\to b_{1}$ and $b_{1}a_{3}\to b_{2}$, we merge them into a single clause $a_{1}a_{2}a_{3}\to b_{2}$.
  2. 2.
    Random drop: Given a clause $a_{1}a_{2}\cdots a_{n}\to b$. We drop each of the antecedents $a\in\{a_{1},a_{2},\cdots a_{n}\}$ at probability $p_{drop}$. We apply this dropping to every clause in the proof except the first one to ensure that we do not drop the premises.
We use $p_{merge}=p_{drop}=p_{skip}$.
Additionally, when flatting the proof trees with topological sort, we break ties randomly. We also randomize the order of the symbols in the antecedents.
[…]
Figure 3: The distribution of lengths and the first step in each paragraph where $z$ is the consequence in the Calcutec dataset. The first/second row are the statistics before/after some steps are randomly dropped.
Appendix G Extra Experiments with Calcutec
G.1 Additional Setups
Unseen Inference Process.
Based on Assumption 2.9 and Assumption 2.5, input-label pairs of a downstream task corresponds to prefix-reference pairs in a paragraph. To examine whether the trained model can generalize well when the induction process for the label is different from the induction process for the pronoun in the training data, we generate a training dataset where all the pronouns are induced from the premise with a left-branching proof tree with a depth equal to 2 (Figure 3(a)), while the test data contains samples whose labels are induced from the input with balanced trees (Figure 3(b)).
Different Input Lengths.
For each downstream tasks, we experiment with examples with different lengths. When the inference process is branching, having input length equal to $4$ makes the proving tree deeper.
No perturbations.
As described in §F, we apply some random perturbations on the proving process. We also experiment with the setup where we do not apply any perturbations.
With/Without Rewriting the First Step.
As described in §E, we rewrite the first step of the proof. We also experiment with the setup where we do not rewrite the first step.
Model Size.
We also experiment with different models sizes. We experiment with GPT-2 models that have 3, 4 and 5 layers.
G.2 Results and Discussion
Unseen Inference Process.
Figure 4(a) and Figure 4(d) show that the ICL performance on the branching examples is similar to the performance on the branching examples. It suggests that the model can generalize to examples that requires an unseen reasoning process. Interestingly, Table 5 show that using chain-of-thoughts mitigates this gap.
Different Input Lengths.
Figure 4(b) and Figure 4(e) show that the model can still do ICL for the examples with length equal to $4$. However, compared with the performance on examples with length equal to $3$ (Figure 4(c) and Figure 4(f)), the performance is worse. This may be because solving these length-4 examples requires more reasoning steps.
With/Without Rewriting the First Step.
Figure 7 shows that models trained with proofs that are rewritten has similar performance as models trained with the proofs that were rewritten (Figure 5). This suggests that rewriting the first step in the proof is not necessary for the model to acquire the ICL ability.
Model Size.
Figure 8 show that deeper models have better ICL performance. It is aligned with the real-world observation that scaling helps model performance.
[…]
(a) The proof tree a paragraph in the training dataset corresponds .
[…]
(b) A balanced tree for a downstream task sample.
Figure 4: Proof trees examples.
Appendix H Hyper-parameters
We train our model using batch size 256, warm up ratio 5%, and we truncate the sequence length to 512 tokens and the default parameters for the optimizer. We use the implementation of GPT-2 by Hugging Face transformers v4.27.2. All models can be trained with 4 RTX 2080ti within 8 hours.
(a) balanced, length = $4$, $y_{1},y_{2}$
(b) branching, length = $4$, $y_{1},y_{2}$
(c) branching, length = $3$, $y_{1},y_{2}$
(d) balanced, length = $3$, $y_{3},y_{4}$
(e) branching, length = $4$, $y_{3},y_{4}$
(f) branching, length = $3$, $y_{3},y_{4}$
Figure 5: In-context learning accuracy with Calcutec when using different verbalizers ($y_{1},y_{2}$ or $y_{3},y_{4}$) and input lengths (3 or 4). The dotted lines represent the performance of unseen combinations described in §4.2, while the different colors represent the number of formulas each class ($v_{+}$ or $v_{-}$) is associated to. The main lines represent the average accuracy of 5 tasks. We plot the performance of each task in lighter colors.
(a) balanced, length = $4$, $r_{1},r_{2}$
(b) branching, length = $4$, $r_{1},r_{2}$
(c) branching, length = $3$, $r_{1},r_{2}$
(d) balanced, length = $4$, $r_{3},r_{4}$
(e) branching, length = $4$, $r_{3},r_{4}$
(f) branching, length = $3$, $r_{3},r_{4}$
Figure 6: In-context learning accuracy with Calcutec when no steps are dropped ($p_{skip}=0$).
(a) balanced, length = $4$, $r_{1},r_{2}$
(b) branching, length = $4$, $r_{1},r_{2}$
(c) branching, length = $3$, $r_{1},r_{2}$
(d) balanced, length = $4$, $r_{3},r_{4}$
(e) branching, length = $4$, $r_{3},r_{4}$
(f) branching, length = $3$, $r_{3},r_{4}$
Figure 7: In-context learning accuracy with Calcutec without rewriting the first step to include contain the premise of the proof.
(a) balanced, length = $4$, 3 layers
(b) balanced, length = $4$, 4 layers
(c) balanced, length = $4$, 5 layers
(d) branching, length = $4$, 3 layers
(e) branching, length = $4$, 4 layers
(f) branching, length = $4$, 5 layers
Figure 8: The in-context learning performance when using models with different model depths.
Branching Balanced
$r_{1},r_{2}$ $r_{3},r_{4}$ $r_{1},r_{2}$ $r_{3},r_{4}$
Task ICL CoT ICL CoT ICL CoT ICL CoT
Single 57.1 91.7 55.6 92.0 68.5 89.8 64.9 90.3
Double 53.5 76.3 51.1 77.1 58.5 76.1 56.2 75.8
Triple 53.0 73.0 51.7 73.4 57.0 68.2 54.2 67.0
Table 5: The 4-shot accuracy of in-context learning (ICL) versus chain-of-thoughts (CoT).
(a) 5-digit
(b) 6-digit
Figure 9: The distribution of the number of reasoning steps in the dataset when some of them are dropped at different probability. Each number is the average over 5 datasets generated with different random seeds.
Appendix I Formal Description of the Digit Addition Data
[…]
(a) 5 digits
(b) 6 digits
Figure 10: The accuracy of the models for the addition tasks. The x-axis represents the probability at which we drop each reasoning step in the training data independently. The solid line represents the ratio of testing samples where the model can output the exact answer, while the dashed line represents the character-level accuracy. The results are the average of 5 random seeds.
Figure 11: The exact accuracy (y-axis, solid points) and digit-level accuracy (y-axis, hollow points) versus validation loss (x-axis) for a 5-digit addition task.
Figure 12: The exact accuracy (y-axis, solid points) and digit-level accuracy (y-axis, hollow points) versus validation loss (x-axis) for a 6-digit addition task.

### v2

@@
-Understanding In-Context Learning with a Pelican Soup Framework
+Pelican Soup Framework: A Theoretical Framework for Language Model Capabilities
 Abstract
-Many existing theoretical analyses of in-context learning for natural language processing are based on latent variable models that leaves gaps between theory and practice.
-We aim to close these gaps by proposing a theoretical framework, the Pelican Soup Framework.
-In this framework, we introduce (1) the notion of a common sense knowledge base, (2) a general formalism for natural language classification tasks, and the notion of (3) meaning association.
-Under this framework, we can establish a $\mathcal{O}(1/T)$ loss bound for in-context learning, where $T$ is the number of example-label pairs in the demonstration.
-Compared with previous works, our bound reflects the effect of the choice of verbalizers and the effect of instruction tuning.
-An additional notion of atom concepts makes our framework possible to explain the generalization to tasks unseen in the language model training data.
-Finally, we propose a toy setup, Calcutec, and a digit addition task that mimics types of distribution shifts a model needs to overcome to perform in-context learning.
-We also experiment with GPT2-Large on real-world NLP tasks.
-Our empirical results demonstrate the efficacy of our framework to explain in-context learning.
-In-Context Learning, Pelican Soup Framework
+In this work, we aim to better understand how pretraining allows LLMs to (1) generalize to unseen instructions and (2) perform in-context learning, even when the verbalizers are irrelevant to the task.
+To this end, we propose a simple theoretical framework, Pelican Soup, based on the logical consistency of the training data, the notion of “reference-meaning association”, and a simple formalism for natural language processing tasks.
+Our framework demonstrates how linguistic and psychology studies can inform our understanding of language models and is connected to several other existing theoretical results.
+As an illustration of the usage of our framework, we derive a bound on in-context learning loss with our framework.
+Finally, we support our framework with empirical experiments and provide possible future research directions.
+Pelican Soup Framework:
+A Theoretical Framework for Language Model Capabilities
+Ting-Rui Chiang University of Southern California tingruic@usc.edu          Dani Yogatama University of Southern California yogatama@usc.edu
 1 Introduction
-Large language models (LM) have demonstrated the ability to perform downstream natural language processing (NLP) classification tasks via in-context learning (ICL) (Chowdhery et al., 2022).
-That is, they can predict the label of an input by conditioning on the demonstration, i.e., a sequence of input-label pairs (Brown et al., 2020), even when the verbalizers (labels present in the demonstration) are semantically irrelevant to the task, e.g., foo/bar instead of negative/positive (Wei et al., 2023).
-To understand why LMs trained with nonstructural general text can perform ICL, a line of works base their theoretical analysis on the assumption that the data is from a latent variable model (Xie et al., 2022; Zhang et al., 2023; Wang et al., 2023b).
-However, the assumptions on a latent variable model in these works leave a gap between theory and practice.
-Xie et al.
-(2022) assumes that the general text for training LMs is from a hidden Markov model, which is an oversimplification of natural language.
-Zhang et al.
-(2023) propose a more general latent variable model for the data generation process.
-Although their model is more realistic than a hidden Markov model, their analyses are based on the assumption that there is a data generation process that generates both the general text for LM training and the demonstration for in-context learning.
-As demonstrations have a format very different from general text, a question to be answered is whether such a generation process really exists.
-Additionally, considering the diversity of the semantics natural language can express, it is likely that the training data does not cover all possible values in the latent variable’s space.
-It requires more characterizations on the latent space to discuss the generalization to semantics unseen in the training data.
-Besides, previous theoretical analyses do not fully answer some questions either (Xie et al., 2022; Zhang et al., 2023; Hahn & Goyal, 2023).
-For example, what is the effect of the choice of verbalizers on ICL performance?
-How does the distribution of downstream task examples affect ICL?
-How instruction-tuning helps ICL (Chung et al., 2022)?
-To answer these questions, we propose a theoretical framework, the Pelican Soup Framework.
-We introduce the notion of a commonsense knowledge base (KB) and define the data generation process as the process that generates documents consistent with the KB.
-We also introduce a formalism for NLP classification tasks.
-This notion and formalism instantiate a data generation process that satisfies the assumption made by Zhang et al.
+Large language models (LLMs) have demonstrated the capability to perform downstream natural language processing (NLP) tasks.
+By following instructions, LLMs can perform tasks with zero-shot examples, demonstrating their reasoning capability.
+With some input-output examples provided in the prompt, LLMs can also perform tasks without instructions, which is known as in-context learning (ICL) (Chowdhery et al., 2022).
+In particular, Brown et al.
+(2020) show that LLMs can perform ICL for classification tasks even when the verbalizers (labels present in the demonstration) are semantically irrelevant to the task, e.g., foo/bar instead of negative/positive Wei et al.
 (2023).
-With slight adaption on the Zhang et al.
-(2023), we establish a similar bound for ICL loss under our framework.
-Additionally, because we see the latent variable as the association between the verbalizers and their meaning, our bound can reflect the effect of the choice of verbalizers.
-It also reflects how instruction tuning affects ICL performance.
-If we further assume that all latent meaning can be represented with logic formula over a finite set of atom concepts as suggested by cognitive science theories Fodor (1975, 2008); Piantadosi (2021), we can characterize the generalization to the unseen meaning during inference time.
-Having this assumption instantiates a setup compatible with the theory for the emergence of complex skills by (Arora & Goyal, 2023), and we can thus apply their analyses on generalization to our framework.
-We also discuss our results with the results by Hahn & Goyal (2023), where they bound ICL loss with the description length of the ground truth input-label mapping function.
-Inevitably, whether our analyses are meaningful depends on how deep models such as transformers (Vaswani et al., 2017) generalize.
-Specifically, as most of the previous theoretical analyses do, we show that a perfect LM that estimates the next token distribution perfectly according to the training data distribution can do ICL.
-However, when doing ICL, LMs are conditioned on the demonstration, which has a low probability mass in the training data.
-LMs can do ICL only if it can generalize well under this distribution shift from training data to ICL demonstrations.
-As the generalization of LMs is difficult to analyze theoretically, we conduct experiments to show empirical evidence for our framework.
-We construct a synthetic dataset, Calcutec (§4), and a synthetic task, the addition task (§5), to mimic the distribution shifts LMs need to overcome.
-With Calcutec, we show that LMs can do ICL regardless of the distributional shifts and observe the efficacy of chain-of-thoughts.
-With the addition task, we observe that larger models generalize better under a certain type of distributional shift, which is aligned with the observation about model scaling in the real world Wei et al.
+However, it is unclear how pretraining leads to these capabilities.
+To explain how LLMs acquire these capabilities, we propose a simple theoretical framework, the Pelican Soup framework in §2.
+Our framework is based on some very general assumptions, such as the logical consistency of training documents (Assumption 3.2) and the variable meanings with which a phrase can be associated in different contexts (Assumption 3.3).
+Our framework also includes a simple formalism for NLP tasks, which helps explain why LMs can follow instructions.
+In §4, we show how we can use this framework to analyze LLM’s ICL capability, which mitigates the limitations of previous theoretical analyses.
+For example, in the first theoretical analyses of ICL, Xie et al.
+(2022) assumes that the general text for training LMs is from a hidden Markov model (HMM), which may be an oversimplification of natural language.
+In comparison, our framework does not require this strong assumption.
+Our framework offers unique insights compared to existing theories.
+For example, although the generation process by Zhang et al.
+(2023) is more general than the HMM assumption by Xie et al.
+(2022), it lacks groundings in real-world linguistic phenomena.
+Our framework mitigates this limitation, as it helps us better explain the physical meaning of the terms in the bound on the ICL loss and shows how the terms reflect real-world practices, such as the choice of verbalizers and the distribution of test examples.
+Compared to the work by Hahn and Goyal (2023), our theory allows us to bound ICL loss without relying assumptions on the grammar that generate the data.
+Furthermore, in §5, inspired by the cognitive science theories Fodor (1975, 2008); Piantadosi (2021), early development of artificial intelligence  (Siskind, 1996; Murphy, 2004) and formal linguistics Carnap and others (1968); Bresnan and Bresnan (1982); Steedman (1987, 1996); Sag et al.
+(1999), we provide an extension of our framework to explain why generalization is possible.
+The extension also connects our framework to other theoretical results.
+For example, our extension instantiates the complex skills in the theory by Arora and Goyal (2023).
+In §6, with an additional assumption, we can relate the ICL bound to the description length of the underlying input-output mapping function as done by Hahn and Goyal (2023).
+Our work informs future LLM research directions.
+Scientifically, we shed light on how properties of the training data contribute to the ICL and the instruction-following capabilities of LLMs.
+The framework also shows how linguistic and psychology studies can inform our understanding of modern NLP.
+Practically, we highlight the importance of acquiring knowledge about the interrelation between concepts through pretraining.
+As shown in previous studies, the language modeling objective is inefficient for knowledge acquisition Allen-Zhu and Li (2023); Chiang et al.
 (2024).
-Finally, in §6, we experiment with natural language and show that using pronouns as verbalizers can lead to comparable performance (with GPT-2 Large), which is aligned with our framework where LMs learn ICL ability by modeling meaning association of words.
+We suggest developing a better pretraining technique is crucial for future NLP development.
+Our proposed experimental setups can also facilitate future studies on LLMs’ acquisition of the capabilities.
 2 The Pelican Soup Framework
-2.1 Motivation: Pelican Soup
+We aim our theoretical framework at explaining why LLMs can perform well on prompts for downstream tasks even though the prompts have a different distribution than the training corpus.
+Therefore, our framework includes assumptions qualifying the training corpus distribution in (§3.1) and a formalism for NLP tasks (§3.2).
+Later, we will show how this framework allows us to bound the loss of ICL.
+3 Motivation
 The Pelican Soup game inspires our framework.
 It is a game involving a puzzle master who has a story in their mind.
-The game participants aim to recover the story by asking the puzzle master yes/no questions.
+The game participants ask the puzzle master questions.
+The goal of the participants is to recover the story based on the answer to the yes/no questions.
 An observation is that once the participants recover the story, they can answer any questions about the story.
-Therefore, the story has a similar role as a latent variable defining the input-output mapping, and the yes/no questions are similar to the demonstrations for in-context learning.
-We include an example in Appendix A.
-Given the above observation, we can study in-context learning by considering why humans can solve Pelican Soup riddles.
-We conjecture that this is because the person who makes the story and the ones who solve the riddle share the same (or similar) commonsense (McCarthy, 1960) about logical relationships among things in this world (Schank & Abelson, 1988).
-This inspires us to introduce the notion of a commonsense knowledge base in our framework.
+This is similar to ICL, where the model is required to “figure out” the underlying task based on the input-output pairs in the given demonstration.
+Once the model identifies the underlying task, it will be able to predict the label of any input.
+Therefore, the story has a similar role as a latent variable defining the task, and the yes/no questions are similar to the demonstrations for ICL.
+We include an example in §A.
+Given the above observation, we can study ICL by considering why humans can solve Pelican Soup riddles.
+We conjecture that this is because the person who makes the story and the ones who solve the riddle share the same (or similar) commonsense (McCarthy, 1960) about logical relationships among the events in this world (Schank and Abelson, 1988).
+Based on this shared commonsense, the game participants can eliminate hypothesis contradictory to the information that the asked yes/no questions reveal, which allows the participants to uncover the underlying story.
+This idea of “shared commonsense” inspires us to introduce the notion of “no ambiguity” and “consistency” in our framework.
 3.1 Training Data Distribution
-Our theory framework is based on the interrelations between the semantics of sentences.
-As suggested by the long history of linguistics studies that aim at representing sentence semantics with formal languages (Carnap et al., 1968; Bresnan & Bresnan, 1982; Steedman, 1987, 1996; Sag et al., 1999), such as first-order logic (Frege et al., 1879; Peirce, 1883), it is reasonable to make a general assumption on language:
-Assumption 2.1 (Sentence).
-We assume that a sentence in a language is a sequence of words such that humans can determine whether one sentence entails another sentence or whether two sentences contradict each other.
-Motivated by the Pelican Soup game and the early formulation of AI (McCarthy, 1960; Croft, 1993), we introduce the notion of a commonsense knowledge base and use it to restrict the possible combination of sentences that can co-occur in a paragraph:
-Assumption 2.2 (Knowledge base).
-There is a knowledge base $\mathrm{KB}$, which is a set of sentences, such that any paragraph with non-zero probability mass is a set of sentences such that for any two sentences $x_{1},x_{2}$ in the paragraph, $x_{1}\wedge\mathrm{KB}$ does not contradict with $x_{2}$.
-It is natural to assume the existence of this knowledge base.
-This knowledge base may include statements describing interrelations among the concepts represented by words, as cognitive psychologists have suggested word semantics are based on these interrelations (Siskind, 1996; Murphy, 2004).
-For example, the knowledge base may contain a rule: “If X moved, then the position of X changed” because the semantic of the verb “move” is embodied in its implication on “position”.
-Based on this knowledge base, humans can do step-by-step reasoning with language.
-Humans also avoid writing paragraphs that include contradictory sentences.
-We also observe that language allows us to arbitrarily associate concepts to an entity or a pronoun, which we refer to as reference.
-For example, when the pronoun “she” or the human name “Emily” is present in a paragraph, it is associated with a person that can be described with a set of sentences.
-The usage of the reference is dependent on the meaning it is associated with and is consistent within its context.
-For example, if “she” is associated with the sentence “a person who has a house”, then the sentence “she has no property” will have 0 probability mass.
-Moreover, when we want to refer to “the person who has a house” instead of repeating the sentence again, we use “she” for an abbreviation (when there is no ambiguity).
-We can see abstract adjectives such as “good” and “bad” as references too, as their meaning also depends on the context.
-However, their meaning may not be as variable as pronouns.
-We reflect this difference with a prior distribution for the meaning a reference is associated with.
-For simplicity, we only consider single-word references and assume that such association is consistent throughout a document. We write down an assumption formally:
-Assumption 2.3 (Reference).
-A set of words $\Gamma\subset\Sigma$ serve as references in the language.
-For every document in the training data, some $r\in\Gamma$ can be associated with a meaning represented as a set of sentences $Z_{r}$, which follows a prior distribution $\Pr(Z_{r})$.
-Any $z\in Z_{r}$, $z$ present in the document can be replaced by $r$.
-Finally, we assume a document is composed of a set of paragraphs where some references are present:
+Our theory framework is based on the interrelations between the semantics of sentences. Thus, we first make a general assumption:
+Assumption 3.1 (No ambiguity).
+For all pairs of sentences $x_{1},x_{2}$ in a language, we assume humans can determine the relationship between $x_{1}$ and $x_{2}$ is contradiction, entailment or neutral.
+Assumption 3.1 allows us to qualify sentences that can cooccur in a paragraph with non-zero probability. That is, a paragraph generally does not contain self-contradictory information:
+Assumption 3.2 (Consistency).
+Any paragraph with non-zero probability mass does not contain two sentences contradicts to each other.
+To show how modeling natural language leads to the ICL capability, we further introduce the notion of expression-meaning association as a latent variable.
+It reflects the fact that language allows us to associate meanings with the surface form expressions quite freely.
+For example, when “she” or the human name “Emily” is present in a paragraph, it is associated with a certain person of certain characteristics, which exhibit its meaning.
+Meanwhile, the usage of the expression depends on the meaning it is associated with and is consistent within its context.
+For example, if “she” is associated with the sentence “a person who has a house”, then by Assumption 3.2, the sentence “she has no property” will have 0 probability mass.
+Moreover, when we want to refer to “the person who has a house” instead of repeating the sentence again, we use “she” as an abbreviation.
+For simplicity, we only consider single-word expressions and assume that such association is consistent throughout a document.
+Assumption 3.3 (Expression-meaning association).
+There is a set of words $\Gamma$ such that for every document in the training data, some $r\in\Gamma$ in the document is associated with a meaning represented as a set of sentences $Z_{r}$ with a prior distribution $\Pr(Z_{r})$.
+Any $z\in Z_{r}$ present in the document can be replaced with $r$ without breaking the logical consistency of the document.
+Adjectives such as “good” and “bad” are also expressions that can be associated with variable meanings, and their meanings also depend on the context.
+However, the association may not be as variable as pronouns’.
+We reflect this with a prior distribution for the meaning with which an expression is associated later in our theoretical analysis.
+Finally, we assume a document is a set of paragraphs where some expressions in $\Gamma$ are present:
 Assumption 3.4 (Document).
-A document is a concatenation of paragraphs separated with a delimiter $d$ (e.g.
-a blank line).
-Also, there is a set of reference $S\subset\Gamma$ such that in each of the paragraphs, a $r\in S$ is present.
-2.3 A Formalism for NLP Classification Tasks
-With the commonsense knowledge base $\mathrm{KB}$ defined in Assumption 2.2, we propose a simple formalism: For any objective or prescriptive NLP classification task (Rottger et al., 2022) that classifies an input $x$ to one of $|\mathcal{Y}|$ classes, it can be described with some instructions $u$ and some descriptions $v_{i}$ for $i=1,2,\cdots,|\mathcal{Y}|$ such that $x$ is labeled as $y_{i}$ if and only if
-$\mathrm{KB}\wedge t\wedge x\models v_{i}.$ (1)
-That is, $x$ belongs to class $y_{i}$ if and only if based on the commonsense rules in $\mathrm{KB}$ and the instruction $t$, $x$ entails $v_{i}$.
-Note that because we can rewrite Eq.
-1 as $\mathrm{KB}\wedge x\models u\to z_{i}$, we can combine the task instructions with the description of each class $v_{i}$ for $i=1,2\cdots,|\mathcal{Y}|$.
-Therefore, we can represent any NLP classification tasks as $\langle v_{1},v_{2},\cdots,v_{|\mathcal{Y}|}\rangle$.
-[…]
-Assumption 2.5 (Downstream Task).
-NLP classification tasks can be defined with the formulation described above.
-Namely, given any classification task with classes $\mathcal{Y}$, we can define it with the descriptions $\{v_{y}\}_{y\in\mathcal{Y}}$.
-We further assume that the descriptions of different classes contradict with each other, so every input belongs to only one class.
-2.4 Bounding ICL Loss
-Our framework is quite general; by adapting and combining the analyses by Zhang et al. (2023) and Hahn & Goyal (2023), we have the following theorem:
-Theorem 2.6 (Average ICL Loss).
-Let the description of a classification task be $\{z_{y}\}_{y\in\mathcal{Y}}$ and $z^{*}$ be the association between the task description $\{v_{y}\}_{y\in\mathcal{Y}}$ and the verbalizers $\Gamma_{\mathcal{Y}}$ used for ICL.
-Let $K$ be the constraints used for decoding, $\dot{g}$ be the event where a document belongs to a certain genera.
-Let $S_{t}=x_{1},r_{2},d,x_{2},r_{2},d\cdots,x_{t},r_{t},d$, where $r_{t}$ is the verbalizer that is associated with the label of $x_{t}$ and $d$ is the delimiter.
-We have for any positive integer $T$, the average cross-entropy loss of ICL is bounded as:
-$\begin{split}&-\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1},K)\\ &\leq\begin{aligned} &-\frac{1}{T}\log\Pr(z^{*},\dot{g}|K)\\ &-\frac{1}{T}\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},\dot{g},S_{t-1},K)\\ &-\frac{1}{T}\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K)}{\Pr(x_{t}|S_{t-1},K)}\end{aligned}\end{split}$ (2)
-When the last two terms on the right-hand side are non-negative, Eq.
-2 shows the average cross-entropy loss of ICL converges to 00 in $\mathcal{O}(1/T)$.
-We discuss the three terms in the right-hand side below.
-The second term becomes 00 if we set $K$ as the constraint that the next two tokens of $S_{t-1},x_{t}$ must be a verbalizer and the delimiter for all $t$.
-This is because Assumption 2.2 ensures that $x_{t}$ does not conflict with $r_{t}$ and Assumption 2.5 ensures that $x_{t}$ conflict with verbalizers other than $r_{t}$.
+A document is a concatenation of paragraphs containing $r\in\Gamma$ separated with a delimiter $d$ (e.g., a blank line).
+Based on the above assumptions, we can define a perfect LM.
+Definition 3.5 (Perfect LM).
+We define an LM to be perfect if it only generates outputs satisfying the above assumptions.
+In the following, we will show that a perfect LM can follow instruction and do ICL.
+3.2 A Formalism for NLP Tasks
+With Assumption 3.1, we propose a simple formalism: For any objective or prescriptive NLP task (Rottger et al., 2022) that maps an input $x$ to a set of acceptable outputs $Y$, that task can be described with some task instructions $u$ such that $u\wedge x\wedge y$ does not cause a contradiction if and only if $y\in Y$.
+For example, if the task is a generation task, such as solving a math word problem, the instruction $u$ may specify the format (e.g., “Think step-by-step and output the answer after ####”) and the input $x$ is a math word problem to be solve.
+LLMs’ response is an acceptable answer if and only if the response does not cause a contradiction to $u$ and $x$, i.e., it follows the instruction and the reasoning process aligns with the problem $x$.
+When it is a classification task, the instruction $u$ should contain descriptions about each label $y$.
+For example, we can formulate the sentiment analysis task over movie reviews as $u=\langle v_{+},v_{-}\rangle=\langle\text{``I like the movie''},\text{``I dislike the movie''}\rangle$.
+Under this formalism, it is trivial that perfect LLMs can follow instructions and solve tasks.
+It is because when the prompt for an LLM is the concatenation of the instruction and the task input, $u;x$, Assumption 3.2 (the consistency assumption) ensures that a perfect LLM only generates $y$ such that $y$ is logically consistent to its prompt $u;x$.
+Based on our NLP task formulation, if $y$ is logically consistent to $u$ and $x$, then $y$ is in $Y$, meaning that $y$ is an acceptable answer.
+Two intricate questions remain, how can LLMs perform ICL? and why is it possible for an LM to generalize to unseen instructions? We discuss these two questions in §4 and §5.
+4 Bounding ICL Loss
+To see how we can analyze ICL with our framework, we first setup a latent variable model with our framework using Assumption 3.3 (the expression-meaning association assumption).
+In this latent variable model, the latent variables are the association between the expressions in the document and their meanings.
+Assumption 3.2 (the consistency assumption) defines the support of the distribution of the latent variable conditioning on a context: Associations that break the logical consistency of the context have zero probability.
+Assumption 3.2 also implies that, give a prefix, continuations that imply invalid associations have zero probability mass.
+We analyzing ICL with this latent variable model by focusing on the latent variable for the association between the verbalizers and the meanings.
+Given a classification task, the underlying latent variable value $z^{*}$ that produces the demonstration is the association where verbalizers are associated with the class descriptions of the task (following the definition of tasks in §3.2).
+For example, when “foo” is used to represent the positive class of an sentiment classification task, “foo” should be associated with the description of the positive class, i.e., “I like the movie”.
+If an LLM can infer this underlying association $z^{*}$ based on the examples in the demostration and output a continuation satisfying Assumption 3.2, then it can predict the correct verbalizer for a test example.
+Formally, adapting and combining the analyses by Zhang et al. (2023) and Hahn and Goyal (2023), we have the following theorem.
+Theorem 4.1 (Average ICL Likelihood).
+Denote a sequence of input-output pairs as $S_{t}=x_{1},r_{1},d,x_{2},r_{2},\cdots,x_{t},r_{t},d$, where $r_{i}$ is the correct verbalizer with which the label of $x_{i}$ is associated for $i=1,2,\cdots,t$ and $d$ is the delimiter that separates the examples.
+Let the description of a task that maps inputs to classes $\mathcal{Y}$ be $\{v_{y}\}_{y\in\mathcal{Y}}$ and $z^{*}$ represents that the task descriptions $\{v_{y}\}_{y\in\mathcal{Y}}$ are associated with the corresponding verbalizers $\{r_{y}\}_{y\in\mathcal{Y}}\subset\Gamma$ used for ICL.
+We have for any $T\in\mathbf{Z}^{+}$, the average log likelihood that the LM predicts the correct verbalizer $r_{t}$ for $t=1,2,\cdots,T$ is bounded as:
+$\begin{split}&\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1})\\ &\geq\begin{aligned} &\frac{1}{T}\log\Pr(z^{*})\\ &+\frac{1}{T}\sum_{t=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},S_{t-1})\\ &+\frac{1}{T}\sum_{t=1}^{T}\log\frac{\Pr(x_{t}|z^{*},S_{t-1})}{\Pr(x_{t}|S_{t-1})}\end{aligned}\end{split}$ (1)
+When the last two terms on the right-hand side are nonnegative, Eq. 1 shows that the average log likelihood converges to 0 in $\mathcal{O}(1/T)$. We discuss the terms on the right-hand side below.
+The second term is 0.
+This is because Assumption 3.2 ensures that the probability $\Pr(r|x_{t})$ is zero if $r$ is not an acceptable output for $x_{t}$ (i.e., when $r\neq r_{t}$ for classification tasks).
 We then look at the last term.
-This term is 00 when $x_{t}$ is conditionally independent to $z^{*}$ as assumed by Zhang et al.
+This term is 0 when $x_{t}$ is conditionally independent of $z^{*}$ as assumed by Zhang et al.
 (2023).
-However, this may be an over-simplification because, in natural language, the transition from $x_{t}$ to its next token depends on the content of $x_{t}$.
-This assumption may actually be unnecessary for convergence because $x_{t}$ is an example from a downstream task related to $z^{*}$; it is likely that
-$\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K)\geq\Pr(x_{t}|S_{t-1},K),$
-which implies that this term is non-negative, and we can thus ignore this term.
-More rigorously, we can show the following corollary:
+However, this may be an oversimplification because, in natural language, the transition from $x_{t}$ to its next token depends on the content of $x_{t}$.
+Fortunately, this assumption may actually be unnecessary for convergence because $x_{t}$ is an example from a downstream task related to $z^{*}$; it is likely that
+$\Pr(x_{t}|z^{*},S_{t-1})\geq\Pr(x_{t}|S_{t-1}),$
+which implies that this term is non-negative, and we can thus ignore this term. More rigorously, we can show the following corollary.
 Corollary 4.2 (Expected Average ICL Loss).
-Let $\dot{g}$ represent a set of documents whose paragraphs are conditionally independent to each other given $z^{*}$, i.e.
-$\Pr(x_{1},d,x_{2},d,\cdots,x_{T},d|z^{*},\dot{g})=\prod_{t=1}^{T}\Pr(x_{t},d|z^{*},\dot{g})$.
-If the downstream task data distribution $\mathcal{D}_{X}$ follow $\Pr(x|z^{*},\dot{g})$, then we can bound the average ICL cross-entropy loss over the downstream task as:
-$\begin{split}&\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\left[-\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1},K)\right]\\ &\leq-\frac{1}{T}\log\Pr(z^{*},\dot{g}|K).\end{split}$ (3)
+Assuming $z^{*}$ is the association suggesting that the sub-sequences separated by the delimiter $d$ are independent of each other, namely, $\Pr(x_{1},r_{1},d,x_{2},r_{2},d,\cdots,x_{t},r_{t},d|z^{*})=\prod_{t=1}^{T}\Pr(x_{t},r_{t},d|z^{*})$.
+If the downstream task data distribution $\mathcal{D}_{X}$ follow $\Pr(x|z^{*})$, then we can bound the average ICL likelihood over the distribution as:
+$\begin{split}&\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\left[\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1})\right]\\ &\geq\frac{1}{T}\log\Pr(z^{*}).\end{split}$ (2)
 The right-hand side of Eq.
 2 characterizes the convergence rate and reflects the difficulty of doing ICL.
-If $z^{*},\dot{g},K$ are independent, then we can see this term is proportional to $\Pr(z^{*})$.
-This implies that when the association between label description and the verbalizer is uncommon in the training data (e.g., associating “positive” to “This movie is bad.”), doing ICL is more difficult.
-Eq.
-2 also allows us to analyze the scenario where we do not constrain the next token of $S_{t-1},x_{t}$ to be a verbalizer while decoding.
-We can replace $\dot{g}$ with $\ddot{g}$ that represents the documents satisfying $\dot{g}$ and the constraint $K$ (i.e., verbalizers always follow $x$).
-[^1] Although $\ddot{g}$ may seem unnatural, this genere of documents corresponds to the PCFG structure assumed in Hahn & Goyal (2023).
-This ensures the second term of Eq.
-2 to be 00.
-The cost of using $\ddot{g}$ makes the first term in bound bigger because $\Pr(\ddot{g}|z^{*})\leq\Pr(\dot{g}|z^{*})$.
-This reflects that doing ICL without constraining the next token is harder.
-$\Pr(\ddot{g}|z^{*})$ may also be related to instruction tuning.
-The training examples for instruction tuning are input-output pairs following some format, such as having an instruction (e.g., “Label the example as positive if …”) at the beginning and a prompt after each input (e.g.
-“[example].
-The sentiment of this comment is”).
-Because for examples that follow the special format, the next token after the prompt is always a verbalizer, these examples belong to the genera $\ddot{g}$.
-Having these examples in the training set would increase $\Pr(\ddot{g}|z^{*})$ and thus make the ICL loss bound converge faster.
-This explains why instruction tuning helps ICL.
-2.5 Extending the Framework with Formal Language
-Assuming a latent model poses a dilemma: Language can encode various meanings, so assuming that the latent space is finite is not reasonable, or the space must be very large.
-However, if the latent space is infinite or is very large, it is possible that the limited training data does not cover the whole space.
+It reflects that when the association between label description and verbalizer is uncommon in the training data (e.g., associating “positive” with “This movie is bad.”), doing ICL is more difficult.
+Note that we can extend the results to generation tasks.
+For generation tasks, we usually use a separator (or a short text span) between $x_{t}$ and $r_{t}$.
+We can see the separator as an expression that can be associated with different meanings, so the latent space for $z$ is the meaning the separator can be associated with and $z^{*}$ is the association between the separator and the task instruction (details in §C).
+5 Generalization
+In §2, we assume a latent model, which, however, poses a dilemma.
+Language can encode various meanings.
+If the sequence length is unconstrained, the corresponding semantic space can even be infinite.
+However, if the latent space is infinite, the limited training data would not cover the entire latent space.
 Without any assumption on the latent space (e.g., the relation between the states in the space), it is impossible to discuss the generalization to unseen latent states.
-Thus, we provide an extension to our framework.
-We can address this dilemma by extending our framework with an assumption stronger than Assumption 2.1:
+Thus, we provide an extension of our theoretical framework.
 Assumption 5.1 (Meaning representation).
-There exists (1) a finite set of atom concepts $\Omega$ and (2) a function $f$ that can map any sentence in language to its meaning represented with a logical formula with operands in $\Omega$ such that for any two sentences $s_{1},s_{2}$, the logical relation between $s_{1}$ and $s_{2}$ is the same as $f(s_{1})$ and $f(s_{2})$.
-This assumption is aligned with the cognitive psychology studies that hypothesize the existence of a set of mental tokens (Fodor, 1975, 2008) and a recent study (Piantadosi, 2021) showing that semantics can be encoded with the combination of only a few symbols.
+There exists (1) a finite set of atom concepts $\Sigma$, (2) a knowledge base $\mathrm{KB}$ consisting of logical rules between the atom concepts in $\Sigma$, and (3) a function $f$ that can map any sentence in language to its meaning represented as a logical formula with operands in $\Sigma$ such that for any two sentences $s_{1},s_{2}$, the logical relation between $s_{1}$ and $s_{2}$ judged by humans is the same as $f(s_{1})$ and $f(s_{2})$ given the rules in the knowledge base $\mathrm{KB}$.
+The three elements of this assumption correspond to theories in various fields.
+The notion of atom concepts aligns with cognitive psychology studies that hypothesize the existence of a set of mental tokens Fodor (1975, 2008) and a recent study Piantadosi (2021) suggesting semantics can be encoded with the combination of only a few symbols.
+The notion of knowledge base follows the early formulation of AI (Siskind, 1996; Murphy, 2004).
+As for the existence of a parsing function $f$, it follows the long history of linguistics studying the relationships between natural and formal languages (Carnap and others, 1968; Bresnan and Bresnan, 1982; Steedman, 1987, 1996; Sag et al., 1999; Frege and others, 1879; Peirce, 1883).
 This assumption suggests that if we have the parsing function $f$, solving NLP tasks only requires a finite-length program that can do logical reasoning by manipulating logical symbols according to logical induction rules.
-If a deep model can learn this program, then it can perform a task even if this task is not in the training data.
-This assumption of a finite $\Omega$ also instantiates the concept of language skills by Arora & Goyal (2023), and their theoretical results are thus applicable.
-2.6 Relating to Function Description Length
-When there is no decoding constrains, $\Pr(r_{t},d|x_{t},z^{*},S_{t-1})$ is related to the difficulty of the example. To see this, we need an additional assumption:
+If a model can learn this program, then it can perform a task even if this task is not in the training data.
+This assumption of a finite $\Sigma$ also instantiates the concept of “language skills” by Arora and Goyal (2023), and their theoretical results are thus applicable.
+Figure 1: Calcutec examples for training, in-context learning (ICL), and chain-of-thought (CoT).
+6 Relating to Description Length
+We can see $\Pr(r_{t},d|x_{t},z^{*},S_{t-1})$ as the difficulty of the example by having an additional assumption:
 Assumption 6.1.
-In some documents in the training data, the paragraphs are constituted with steps in a logical induction process with some steps randomly dropped.
-This kind of document may be pervasive in the training data.
+In some documents in the training data, the paragraphs are constituted with steps in a logical induction process, with some steps randomly dropped.
+This kind of document may be prevalent in training data.
 Essays arguing some claims are one example.
@@
 $\Pr(r_{t},d|x_{t},z^{*},S_{t-1})\leq c\cdot\ell(x_{t}),$ (3)
-where $\ell(x_{t})$ is the number of reasoning steps required to solve the task, and $c$ is a constant This $\ell(x_{t})$ corresponds to the description length of the function that maps the inputs to their label in the loss bound by Hahn & Goyal (2023) (more discussion in Appendix D).
-Train
-x57 x56 x64 r3 $\to$ x79 , r1 x57 $\to$ x58 , x90 x58 $\to$ r3 , r3 r1 $\to$ x20 , … ; x80 x66 x63 x83 x1 $\to$ x82 , x80 x82 $\to$ r1 , … , x64 x80 $\to$ x54 .
-ICL
-x44 x67 x34 x62 $\to$ r2 ; x55 x38 x50 x48 $\to$ r1 ; x21 x59 x57 x86 $\to$ r2 ; x55 x76 x84 x99 $\to$
-CoT
-x44 x67 x34 x62 $\to$ x16 , x34 x62 $\to$ x99 , x99 x16 $\to$ r1 ; x77 x34 x62 x97 $\to$ x12 … ; x21 x59 x57 x86 $\to$ x69 , x59 x57 $\to$ x75 , x69 x75 $\to$ r2 ; x55 x76 x84 x99 $\to$
-Table 1: Calcutec examples for training, in-context learning (ICL), and chain-of-thought (CoT).
-3 Distribution Shifts from Language Modeling to In-context Learning
-In addition to the theoretical analyses, an important aspect of understanding ICL is to understand how models generalize.
-The bounds we have established in § 2.4 are based on the analyses of the marginal distribution conditioned on the input $S_{t-1},x_{t}$.
-Because the distribution of prefix $S_{t-1},x_{t}$ is different from the distribution of general text, a model needs to overcome several types of distribution shifts to predict the marginal distribution precisely.
-We characterize a few of them below.
-Format mismatch.
-The demonstration is in the format $x_{t},r_{t},d$, where $x_{t}$ follows the downstream task distribution.
-However, in the LM training data, sequences in this format are unlikely to happen frequently.
-Verbalizer Mismatch.
-The verbalizers appear very frequently in $S_{t-1}$. However, the verbalizer may not have as high frequency in the training data distribution.
-Unseen Reference-Meaning Association.
-When performing ICL, the verbalizers are associated with descriptions of the classes for the task. However, the training data may not include this kind of association.
-We empirically investigate these types of distribution shifts in the next section.
-4 Experimenting with Synthetic Datasets
-As a concrete instance of our framework, we present a toy setup, Calcutec, which includes a data generation process for LM training data and a process for downstream tasks.
-Calcutec follows our framework and mimics the distribution shifts LMs need to overcome as discussed in §3.
-We will use this setup to inspect models’ generalization ability.
-4.1 Calcutec
-Setup.
-Following the assumption in §2, we construct a pseudo-language with the following components:
-  • Logic model: We use a subset of propositional logic as our logic model.
-We only consider Horn clauses (Horn, 1951), i.e., formulas in the form $A\wedge B\to C$.
-This allows us to efficiently generate proof with time complexity linear to the size of the knowledge base (Dowling & Gallier, 1984).
-  • Atom concepts: We use 100 variables as our set of atom concepts $\Sigma$.
-  • $\mathrm{KB}$: For each variable $\sigma\in\Sigma$, there are 5 formulas in the knowledge base whose consequence (the literals after $\to$ in a horn clause) is $\sigma$, while the antecedents (the literals before $\to$) are sampled from $\Sigma\backslash\{\sigma\}$ uniformly at random.
-  • We have a set $\Gamma=\{r_{i}\}_{i=1}^{4}$ representing the reference words described in Assumption 2.3.
+where $\ell(x_{t})$ is the number of reasoning steps required to solve the task, and $c$ is a constant.
+This $\ell(x_{t})$ corresponds to the description length of the function that maps the inputs to their label in the loss bound by Hahn and Goyal (2023) (more discussion in Appendix E).
+7 Inspecting Generalization Empirically
+Although we have shown that perfect LMs can follow instructions (§3.2) and perform ICL (§4) when training data satisfy the assumptions in §2, it remains unverified whether imperfect real-world LMs exhibit the same capabilities.
+For LMs to demonstrate these abilities, they must generalize well from the training set to the prompts used during inference.
+However, analyzing how deep models generalize inherently requires additional assumptions about the model, which is beyond the scope of this paper.
+Thus, to address this gap, we leverage our extended framework (§5) to characterize distribution shifts and design experiments that empirically examine the generalization of LM.
+7.1 Inspecting the ICL Capability
+Firstly, we present a synthetic setup, Calcutec, as a concrete instantiation of our theoretical framework.
+With Calcutec, we show that Transformers can acquire ICL capability by modeling the linguistic characteristics specified in our framework.
+7.1.1 Calcutec
+Setup
+Following the extended framework in §5, we construct a pseudo-language:
+  • Logic model: We use a subset of propositional logic as our logic model. We only consider Horn clauses (Horn, 1951), i.e., formulas in the form $A\wedge B\to C$.
+  • Atom concepts: We use 100 symbols as our set of atom concepts $\Sigma$.
+[…]
+  • We have a set $\Gamma=\{r_{i}\}_{i=1}^{4}$ representing the expressions described in Assumption 3.3.
 Training Dataset.
@@
 In our synthetic language model training dataset, each document contains 16 paragraphs.
-Before starting the generation process, we randomly draw a reference symbol $r_{a}\in\Gamma$ and associate it with two formulas $s_{1},s_{2}$.
-For simplicity, we make these two formulas single-variable, i.e., $s_{1},s_{2}\in\Sigma$.
-Because paragraphs in the real world are sentences ordered in a certain way, we follow Assumption 2.9 and randomly generate a proof as a paragraph.
-Each paragraph represents the induction process of $P\models g$ for some randomly selected $P\subset\Sigma$ and $g\in\Sigma$ that allows us to use the reference symbol $r$ in the paragraph.
-Each sentence in the paragraph is a sentence representing a reasoning step.
-Because we use Horn logic as our logic model, a step is a rule in $\mathrm{KB}$ in the form $a_{1}a_{2}\to b$.
-We separate the clauses in the sequence with commas.
-To simulate the fact that documents in the real world always skip some reasoning steps, we further apply some perturbations over the generated paragraphs that drop some reasoning steps with a skip rate $p_{skip}$ (details in Appendix E).
-[^2] Models can acquire in-context learning ability even with $p_{skip}=0$ (Figure 6 in thr appendix).
-After we generate a document, we replace $s_{1}$ and $s_{2}$ with the reference symbol $r_{a}$.
-We randomly draw another reference symbol $r_{b}\in\Gamma$ and replace the two most frequent symbols in the document with $r_{b}$.
-(We provide the pseudo-code Alg.
-1 in the appendix.)
+Following Assumption 3.2, each paragraph consists of sentences logically coherent to each other.
+Because sentences in the real world are not ordered arbitrarily, we follow Assumption 6.1 and generate random paragraphs following the structure of logical proofs.
+Each paragraph represents an induction process of $P\models g$ for some randomly selected $P\subset\Sigma$ and $g\in\Sigma$.
+Each sentence in the paragraph is a sentence representing a reasoning step for $P\models g$.
+We separate the sentences in the sequence by commas.
+To simulate the noise in the real world, we further apply perturbations that skip some steps with a skip rate $p_{skip}$.
+[^1] Models can acquire in-context learning ability even with $p_{skip}=0$ (Figure 6 in the appendix).
+Following Assumption 3.3, after we generate a document, we randomly associate some symbols $A,B\subset\Sigma$ with $r_{a},r_{b}\in\Gamma$ respectively.
+We reflect this association in the generated document by replacing symbols in AA and $B$ with expression $r_{a},r_{b}\in\Gamma$ respectively (details in Appendix F.1).
 Downstream Tasks.
-Following the formulation in Assumption 2.5, we define a binary classification task by defining the description $v_{+}$ and $v_{-}$ for the positive and negative classes, respectively.
-$v_{+}$ and $v_{-}$ are the disjunctions of atom concepts, i.e., in the form of $a_{1}\vee a_{2}\vee\cdots$.
+Following the formalism in §3.2, we define a binary classification task by defining the descriptions $v_{+}$ and $v_{-}$ of the positive and negative classes, respectively.
+We use the disjunctions of atom concepts (i.e.
+in the form of $a_{1}\vee a_{2}\vee\cdots$) as descriptions of classes.
 We create five downstream tasks using different disjunctions.
-Each input is a subset of variables in $\Sigma$ from which we ensure that only one of the classes can be induced.
+Each input is a subset of variables in $\Sigma$ from which we ensure only one of the classes can be induced.
 Demonstration.
-We represent an input-label pair as $x_{1}x_{2}\cdots\to r$, where $x_{1}x_{2}\cdots$ is the input part and $r\in\{r_{+},r_{-}\}\subset\Gamma$ is a reference symbol serving as a verbalizer.
+We represent an input-label pair as $x_{1}x_{2}\cdots\to r$, where $x_{1}x_{2}\cdots$ is the input part and $r\in\{r_{+},r_{-}\}\subset\Gamma$ is an expression in $\Gamma$ serving as the verbalizer.
 Chain-of-thought.
-A chain-of-thought is in the format same format as the training data, but ends with a reference symbol $r\in\{r_{+},r_{-}\}$, e.g., $x_{1}x_{2}\cdots\to x_{3};x_{3}\cdots x_{4}\to r_{+}$.
+A chain-of-thought is in the same format as the training data but ends with an expression $r\in\{r_{+},r_{-}\}$, e.g., $x_{1}x_{2}\cdots\to x_{3};x_{3}\cdots x_{4}\to r_{+}$.
 This chain-of-thought reflects the step-by-step induction process from the inputs to the label.
-4.2 Inspecting the Distribution Shifts
-Our Calcutec dataset allows us to inspect some types of distribution shifts in § 3.
+We show an example in Figure 1.
+7.1.2 Distribution Shifts
+We make experimental designs to simulate the real-world distribution shifts from training to inference:
 Format Mismatch.
-The training data contains reasoning steps but these steps are absent in the demonstration.
+The reasoning steps are present in the training data but not in the prompts.
 Verbalizer Mismatch.
-When we are picking the reference symbols in $\Gamma$, we assign probability mass 45%, 45%, 5%, 5% to $r_{1},r_{2},r_{3},r_{4}$.
-In this way, we can inspect whether using less frequent verbalizers will lead to worse performance.
+When we choose the expressions in $\Gamma$, we assign the probability mass 45%, 45%, 5%, 5% to $r_{1},r_{2},r_{3},r_{4}$. In this way, we can inspect the effect of using less frequent verbalizers.
 Unseen Tasks.
-To investigate whether the model can generalize to a new combination of formulas unseen in the training data when we generate our training data, we ensure that the references are never associated with some combinations of atom concepts ($s_{1},s_{2}\in\Sigma$).
-We then test the trained model on tasks where $v_{+}$ and $v_{-}$ are the disjunctions of unseen combinations.
-Additionally, while in the training data, each reference symbol is associated with only two atom concepts; we test whether the model can perform in-context learning on tasks where $v_{+}$ and $v_{-}$ are the disjunctions of three atom concepts.
-4.3 Experiment Details
-We use $p_{skip}=0.25$ in our experiment.
-We generate 60,000 documents with 16 paragraphs following the process described in § 4.
+To investigate whether the model can generalize to a new combination of formulas unseen in the training data when we generate our training data, we ensure that the expressions in $\Gamma$ are only associated with the disjunctions of two atom concepts $s_{1},s_{2}$ from a strict subset of all possible combinations $\Sigma\times\Sigma$.
+We then test the trained model on tasks where $v_{+}$ and $v_{-}$ are the disjunctions of the unseen combinations.
+We also test the models on tasks where $v_{+}$ and $v_{-}$ are the disjunctions of three atom concepts $\in\Sigma\times\Sigma\times\Sigma$.
+7.1.3 Experiment Details
+We use $p_{skip}=0.25$ in our experiment, generating 60,000 documents with 16 paragraphs (as described above).
 Among them, we use 10k for validation.
-We train a 6-layer Transformer (Vaswani et al., 2017) model until the loss on the validation set converges.
-We include additional setups in §G
-4.4 Experiment Results and Discussion
-(a) Verbalizers = $r_{1},r_{2}$
-(b) Verbalizers = $r_{3},r_{4}$
-Figure 1: In-context learning accuracy with Calcutec when using different verbalizers ($r_{1},r_{2}$ or $r_{3},r_{4}$).
-The dotted lines represent the performance of unseen combinations described in §4.2.
-The colors represent the number of formulas in $\Sigma$ each class ($v_{+}$ or $v_{-}$) is associated with.
-The main lines represent the average accuracy of 5 tasks.
-We plot the performance of each task in lighter colors.
+We train a 6-layer Transformer (Vaswani et al., 2017) model until the loss on the validation set converges using the standard autoregressive loss.
+We include additional setups in §F.3.
+7.1.4 Results and Discussion
 $r_{1},r_{2}$ $r_{3},r_{4}$
@@
 Triple 53.0 73.0 51.7 73.4
-Table 2: The 4-shot accuracy of in-context learning (ICL) versus chain-of-thought (CoT) using different verbalizers.
-task direct pronoun
-SST-2 63.0 65.3
-CR 61.7 62.9
-MR 59.2 56.7
-Subj 51.0 62.2
+Table 1: The 4-shot accuracy of ICL versus chain-of-thought (CoT) using different verbalizers.
+Figure 2 shows that LMs trained with Calcutec can perform in-context learning.
+This evidence supports our Pelican Soup framework and aligns with our theoretical analysis in §4: The ICL accuracy follows a trend $\exp(c/T)$ for some $c<0$, as our theoretical result in Eq.
+1.
+We further inspect the ICL performance under the distribution shifts described in §7.1.2.
+For the infrequent verbalizer, we observe similar performance regardless of the frequency of the verbalizers ($r_{1},r_{2}$ versus $r_{3},r_{4}$).
+For the unseen tasks, Figure 2 shows that the model has similar performance in tasks defined with seen and unseen combinations of atom concepts (dotted and solid lines), even generalizing to tasks defined with three latent concepts (green lines).
+In sum, the results show that the model can generalize well under several distributional shifts.
+(a) Verbalizers = $r_{1},r_{2}$
+(b) Verbalizers = $r_{3},r_{4}$
+Figure 2: In-context learning accuracy with Calcutec when using different verbalizers ($r_{1},r_{2}$ or $r_{3},r_{4}$).
+The dotted lines represent the performance of unseen combinations described in §7.1.2.
+The colors represent the number of atom concepts each class ($v_{+}$ or $v_{-}$) is associated with.
+The main lines represent the average accuracy of 5 tasks.
+Lines in the lighter color represent the individual tasks.
+We also experiment with 4-shot learning using chain-of-thought.
+Table 1 shows that the model also benefits from chain-of-thought.
+We conjecture that it is because chain-of-thought has a format more similar to the format for training.
+7.2 Inspect Instruction-following Capability
+In addition to ICL, we inspect another intriguing capability of LLMs, the capability of following unseen instructions.
+Our extended framework in §5 suggests that we can see following unseen instructions as doing a sentence completion task that involves atom concepts (or say “skills” in Arora and Goyal (2023)) whose composition is different from the ones involved when completing sentences in training data.
+We also suggest that LMs acquire the capability of following these unseen instructions by modeling the interrelations between the atom concepts involved.
+To show the plausibility of this, we experiment with an arithmetic task.
+7.2.1 Arithmetic Task
+Single-step 0 + 1 * 2 + 3 * 4 = 10 .
+Multi-step 0 + 1 * 2 = 2 . 3 * 4 = 12 . 2 + 12 = 14 .
+Test - seen 2 * 3 + 4 * 5 + 6 =
+Test - unseen 2 + 10 + 4 * 11 * 6 =
+Table 2: Training (first two rows) and test (last two rows) examples used in the arithmetic task §7.2.1.
+We utilize the algebraic structure of the integers under modulo addition/multiplication to construct a language where each expression is constantly associated with a meaning and the atom concepts include 0 to 15, the symbol “+”, “*”, “=”, “.”.
+Each sentence in this language is in the form of an equation with additions and multiplications.
+We generate LM training data in a multi-step and a single-step setup.
+In the multi-step setup, each example contains the steps required to compute the result, each of which involves two to five numbers (details in §G).
+We expect that LMs can learn the interrelation between atom concepts by modeling these steps.
+In the single-step setup, the final result is computed in a single step.
+To simulate the scenario where completing the prompts in the test set involves a different distribution of atom concepts from that of the training examples, we ensure that the five numbers for each training example are either all from $\{0,\cdots,9\}$ or all from $\{6,\cdots 15\}$.
+We then evaluate the model with a seen and an unseen setups.
+In the seen setup, we prompt the model with 5 numbers sampled in the same way as the training set.
+In the unseen setup, we prompt the model to complete equations with 5 numbers where the first, third, and fifth numbers are from $\{0,\cdots,5\}$ and the other two numbers are from $\{10,\cdots,15\}$ (examples in Table 2).
+7.2.2 Results and Discussion
+test $\backslash$ train Multi-step Single-step
+Seen 97.5 (0.3) 99.3 (0.6)
+Unseen 75.3 (2.0) 49.7 (1.0)
+Table 3: The average accuracy of LMs trained for the arithmetic tasks with 5 random seeds.
+Table 3 shows that all LMs achieve a near-perfect accuracy in the seen setup but the LMs trained with the multi-step training set achieve significantly better accuracy in the unseen setup.
+This indicates that modeling the multiple steps in the training data may allow the model to learn the interrelation between the atom concepts and generalize to prompts that involve unseen compositions of atom concepts to some extent, as suggested by our extended framework in §5.
+This is also aligned with the success of symbolic chain-of-thought distillation (Li et al., 2023a; Hsieh et al., 2023; Shridhar et al., 2023).
+task SST-2 CR MR Subj
+direct 63.0 61.7 59.2 51.0
+direct w/ foo/bar 51.7 53.7 52.2 57.1
+pronoun 65.3 62.9 56.7 62.2
 Table 4: The accuracy of using task-specific templates/verbalizers (direct) (Min et al., 2022a) v.s. using task-agnostic templates/pronouns for 16-shot in-context learning with GPT2-Large.
-Training
-4 9 2 8 4 6 + 0 8 0 3 5 0 = 0 0 0 0 0 0 ; 0 9 2 8 4 6 + 0 8 0 3 5 0 = 4 0 0 0 0 0 ; 0 0 2 8 4 6 + 0 0 0 3 5 0 = 4 7 1 0 0 0 ; 0 0 0 8 4 6 + 0 0 0 3 5 0 = 4 7 3 0 0 0 ; $\cdots$ 0 0 0 0 0 0 + 0 0 0 0 0 0 = 4 7 3 1 0 7 ;
-Testing
-8 7 4 0 1 6 + 0 9 2 1 5 0 = 0 0 0 0 0 0 ; 0 0 0 0 0 0 + 0 0 0 0 0 0 =
-Table 4: Training and testing examples for the digit addition task.
-Figure 1 shows that the model trained with Calcutec can perform in-context learning, indicating that the model can generalize under the distribution shifts described in §3.
-This evidence supports our Pelican Soup hypothesis.
-We further inspect the in-context learning performance under the distribution shifts described in §4.2:
-  • Infrequent verbalizer: We observe similar performance of using frequent reference symbols ($r_{1},r_{2}$) or infrequent reference symbols ($r_{3},r_{4}$) as the verbalizers.
-  • Unseen tasks: Figure 1 shows that the model has similar performance over tasks defined with unseen and unseen combinations of atom concepts (dot lines and solid lines).
-The models can also generalize to tasks defined with three formulas (green lines).
-In sum, the results show that the model can generalize well under several distributional shifts.
-We experiment with 4-shot learning using chain-of-thought.
-Table 2 shows that the model also benefits from chain-of-thought.
-We conjecture that it is because chain-of-thought has a format more similar to the format of training.
-5 Digit Addition Task
-In § 4, we show that transformer models can generalize to downstream tasks where the intermediate reasoning steps are missing.
-As this is related to models’ ability to predict $P(r|x,K)$2.6) when we force the model to skip the intermediate steps, we are curious whether transformer models can still generalize even when more steps are missing.
-Since in the Calcutec environment, it is difficult to construct a task that requires a long reasoning process without exponentially increasing the size of the atom concept set $\Sigma$, We propose another toy task, the digit addition task, which involves longer reasoning processes.
-5.1 Setup
-The digit addition task requires the model to perform the addition operation over two $n$-digit numbers.
-It requires $n$ steps to solve if we solve it digit-by-digit.
-We generate a training set consisting of digit-by-digit reasoning steps for 100,000 pairs of numbers.
-Each intermediate step may be dropped at a constant probability $p_{drop}$.
-After training a Transformer model with the language modeling objective, we test whether the model can generate the final answer without generating the intermediate steps.
-We experiment with different models of different sizes.
-From this, we can verify whether models of different sizes can acquire the intuition from the language modeling task.
-We show examples of our data in Table 4.
-Each digit sequence represents a number from the lower digit to the higher digit.
-The reasoning process in the training set gradually updates both sides of “=” from the lowest digit to the highest digit.
-As for the testing example, we skip the intermediate steps, making the model complete right-hand side of “=” in the last step.
-The model needs to acquire intuition from the training set in order to solve the testing samples.
-(We include a rigorous description in Appendix I.)
-Figure 2: The exact accuracy (y-axis, solid points) and digit-level accuracy (y-axis, hollow points) versus validation loss (x-axis) for the 5-digit addition task for five random seeds and different dropping rates $p_{drop}$ (In the appendix, we include the complete results in Figure 11 and the results of the 6-digit addition task in Figure 12.)
-5.2 Results and Discussion
-We report the exact match accuracy and the digit-level accuracy of models trained with different $p_{drop}$ in Figure 2 with 5 random seeds.
-A higher accuracy implies the better intuition the model acquires from the sequence modeling task.
-The results show that three of the four models can achieve perfect accuracy when $p_{drop}$ is as small as $0.3$.
-It suggests that it is possible that the model can predict the final answer even without intermediate steps, even when examples in the training data are mostly step-by-step.
-We also observe that larger models tend to have higher and more stable accuracy.
-When the number of digits is 6 (Figure 12 in the appendix), only the largest model can achieve perfect accuracy.
-This observation is aligned with the emergence of large language models’ ability.
 8 Real-world Evidence
-We inspect whether LMs can do ICL with pronouns well because pronouns are reference words frequently associated with different meaning and our framework suggests that LMs learn ICL ability by modeling the association between reference words and their meaning.
-We thus experiment with the template [input]”, [verbalizer] thought.
-and use “he”, “she” as the verbalizers.
+We inspect how likely that LMs learn ICL by modeling the expression-meaning association as our framework suggests.
+We specifically inspect a small model GPT2-Large and the setting in which we use pronouns as verbalizers.
+Because pronouns are reference words frequently associated with different meanings, we expect that even a small model can do ICL well with pronouns.
+To do so, we experiment with the template [input]”, [verbalizer] thought.
+and use “he”, “she” as verbalizers.
 We follow the setup in Min et al.
-(2022a) and compare the accuracy of the binary classification tasks using GPT2-Large.
+(2022a) and compare the accuracy of binary classification tasks, including SST-2 Socher et al.
+(2013), CR Hu and Liu (2004), MR Pang and Lee (2005), and Subj Pang and Lee (2004), using GPT2-Large.
 Table 4 shows that this task-agnostic template with pronouns is competitive with those task-specific templates.
-This contradicts with the belief that only larger models can do in-context learning with task-irrelevant verbalizer Wei et al.
-(2023).
-It suggests that modeling reference-meaning may indeed contribute LMs’ ICL ability.
+This aligns with our conjecture that the high frequency of pronouns in the training set allows smaller LMs to acquire the ICL capability.
+It also shows that, unlike what Wei et al.
+(2023) claims, not only larger models can do in-context learning with task-irrelevant verbalizers.
+On the other hand, using task-irrelevant verbalizers “foo” and “bar” has lower performance (which aligns with the observation of Wei et al.
+(2023)).
+This may be because learning to perform ICL with those low-frequency words requires more training data, and thus a larger model size.
 9 Related Work
@@
 (2022) suggested that in-context learning is an implicit Bayesian inference process.
-Hahn & Goyal (2023) defined the generation process with Compositional Attribute Grammar, which is weaker than the HMM assumption, explaining the in-context learning ability with the minimum description length.
+Hahn and Goyal (2023) defined the generation process with Compositional Attribute Grammar, which is weaker than the HMM assumption, explaining the in-context learning ability with the minimum description length.
+They also studied the compositionality of natural language tasks with function compositions.
 Zhang et al.
 (2023) assumed a more general latent variable model.
-Arora & Goyal (2023) analyze the emergence of skills based on the scaling law (Hoffmann et al., 2022).
-While their analysis assumes a set of skills as the atomic elements of NLP tasks, our framework is based on a set of atom concepts.
+Arora and Goyal (2023) analyze the emergence of skills based on the scaling law (Hoffmann et al., 2022).
+While their analysis assumes a set of atomic skills for NLP tasks, our framework is based on a set of atom concepts.
 There were also many empirical studies on the in-context learning ability.
-Some works focused on the effect of the instruction (Webson & Pavlick, 2022; Lampinen et al., 2022; Jang et al., 2023), while some focused on the examples in the demonstration (Liu et al., 2022; Lu et al., 2022; Sorensen et al., 2022; Min et al., 2022b; Yoo et al., 2022; Ye et al., 2023; Chang & Jia, 2023; Ye et al., 2023; Wang et al., 2023b; Kossen et al., 2023).
+Some works focused on the effect of the instruction (Webson and Pavlick, 2022; Lampinen et al., 2022; Jang et al., 2023), while some focused on the examples in the demonstration (Liu et al., 2022; Lu et al., 2022; Sorensen et al., 2022; Min et al., 2022b; Yoo et al., 2022; Ye et al., 2023; Chang and Jia, 2023; Ye et al., 2023; Wang et al., 2023b; Kossen et al., 2023).
 Shin et al.
 (2022) found that not all training corpora led to in-context learning ability.
-Prystawski & Goodman (2023) used synthetic data to suggest that the pretraining dataset’s locality structure contributes to the reasoning steps’ effectiveness.
+Prystawski and Goodman (2023) used synthetic data to suggest that the pretraining dataset’s locality structure contributes to the reasoning steps’ effectiveness.
 Wang et al.
@@
 Li et al.
-(2023) provided generalization bounds based on the stability of Transformer models and the distance of downstream tasks.
-Compared to these works, we focus on how the pretraining data in natural language contributes to the in-context learning ability.
-8 Conclusion
-In this work, we propose a framework that instantiates a latent variable model for analyzing ICL.
-Compared with previous works (Xie et al., 2022; Zhang et al., 2023), our latent model better reflects the complexity of language.
-Under our framework, we establish an ICL loss bond that provides insights into the effect of the choice of verbalizers, downstream task distribution, and instruction tuning.
-We further provide extensions to the framework that allow us to analyze the generalization to unseen tasks using the results by Arora & Goyal (2023) and relate our bound to the function description length discussed by Hahn & Goyal (2023).
-We then present a toy setup, Calcutec, as a concrete instance of our framework.
-Calcutec mimics the distribution shifts LMs need to overcome to perform ICL.
-Our empirical results show that transformers trained with autoregressive loss can generalize under these distributional shifts.
-We additionally use an addition task to show that larger models generalize better, potentially explaining why scaling up model size helps LM performance.
-Finally, we experiment with real-world NLP tasks and show that GPT-Large can do ICL with pronouns, providing evidence supporting our framework.
-We expect that our toy datasets can facilitate future research on ICL.
-Impact Statements
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Acknowledgements
-We appreciate Joshua Robinson, Ollie Liu, and Xinyan Velocity Yu for providing valuable feedback on this paper.
+(2023b) provided generalization bounds based on the stability of Transformer models and the distance of downstream tasks.
+We instead focus on how the pretraining data in natural language contributes to the ICL learning ability.
+10 Conclusion and Future Work
+In this work, we propose a framework that explains how linguistic phenomena in the training corpus lead to LLMs’ ICL and instruction-following capability.
+Compared with previous works Xie et al.
+(2022); Zhang et al.
+(2023), our latent model better reflects the complexity of language.
+By introducing the notion of knowledge base and logic system, our framework provides insights into how LLMs can generalize from pretraining to downstream tasks, instantiating a setup compatible with the assumptions made by Arora and Goyal (2023).
+We also relate our bound to the function description length discussed by Hahn and Goyal (2023).
+Our framework illuminates a few possible directions for improving LLMs:
+  1. 1.
+    Our work highlights the importance of learning the interrelation between meanings.
+As previous works have shown that the language modeling objective is inefficient for this purpose (Allen-Zhu and Li, 2023; Chiang et al., 2024), we suggest that developing a more sophisticated learning algorithm is crucial.
+  2. 2.
+    Our theory illustrates how linguistic studies on parsing functions can inform LLM research, offering a new theoretical foundation for analyzing LLM generalization.
+  3. 3.
+    The experimental results of our arithmetic task show that Transformer models can generalize to unseen prompts by modeling the intermediate step-by-step reasoning process.
+This may be related to the success of the symbolic chain-of-thought distillation (Li et al., 2023a; Hsieh et al., 2023; Shridhar et al., 2023).
+Investigating and strengthening the mechanism can improve the efficiency of LM training.
+11 Limitations
+A limitation of our framework is that, as most theoretical studies do, we simplify the real-world scenario to draw insights.
+One simplification we make is that, we do not take the noise in LLMs’ training data into account.
+As we may need to make more assumption on the noise to establish a generic theoretical result, we leave it for future study.
+Another simplification is that, we assume that the language model can perfectly model the distribution of natural language (i.e., only generates outputs satisfying the assumptions in §3.1, especially Assumption 3.2).
+However, it is unlikely to be the case in practice.
+On the one hand, the training data may not cover all the test cases.
+On the other hand, LLMs may not perfectly generalize from the training set.
+We need to make more assumptions on the training/test data distribution and/or have a deeper understanding on how deep learning models generalize to alleviate this assumption.
+Therefore, we deem this out of the scope of this paper.
+We also note some limitations of our experiments.
+As commonly seen in theoretical work, our experiment setup in §7.1 is a simplification of natural language.
+Our intention is to isolate and illustrate specific theoretical mechanisms without confounding variables introduced by the full complexity of natural language.
+The experiment in §8 shows the correlation between better ICL performance and the use of pronouns as verbalizer.
+Although this serves as evidence aligning with our theoretical framework, more studies are required to establish the causal relationship.
 Appendix A A Pelican Soup Example
+We include an example of a Pelican Soup game:
 Puzzle master: A men walks into a restaurant and orders pelican soup. After taking a sip, he loses his mind. Why?
@@
 Appendix B Proof of Theorem 1
-Let $S_{t}=x_{1},r_{2},x_{2},r_{2},\cdots,x_{t},r_{t}$.
-$\Pr(z,g|S_{t},K)=$ $\frac{\Pr(S_{t}|z,g,K)\Pr(z|K)}{\sum_{z}\Pr(S_{t}|z,g,K)\Pr(z,g|K)}$
-$=$ $\frac{\Pr(z^{\prime},g|K)\prod_{i=1}^{t}\Pr(x_{t},r_{t},d|z,g,S_{t-1},K)}{\sum_{z^{\prime},g}\Pr(z,g|K)\prod{i=1}^{t}\Pr(x_{t},r_{t},d|z^{\prime},g,S_{t-1},K)}$
-$P(x_{t+1},r_{t+1},d|S_{t},K)=$ $\sum_{z,g}\Pr(x_{t+1},r_{t+1},d|z,S_{t},K)\Pr(z,g|S_{t},K)$
-$=$ $\frac{\sum_{z,g}\Pr(z,g|K)\prod_{i=1}^{t+1}\Pr(x_{t},r_{t},d|z,g,S_{t-1},K)}{\sum_{z^{\prime}}\Pr(z,g|K)\prod_{i=1}^{t}\Pr(x_{t},r_{t},d|z^{\prime},g,S_{t-1},K)}.$
+Let $S_{t}=x_{1},r_{2},d,x_{2},r_{2},d\cdots,x_{t},r_{t},d$.
+$\Pr(z|S_{t})$
+$=$ $\frac{\Pr(S_{t}|z)\Pr(z)}{\sum_{z}\Pr(S_{t}|z)\Pr(z)}$
+$=$ $\frac{\Pr(z)\prod_{i=1}^{t}\Pr(x_{i},r_{i},d|z,S_{i-1})}{\sum_{z^{\prime}}\Pr(z^{\prime})\prod_{i=1}^{t}\Pr(x_{i},r_{i},d|z^{\prime},S_{i-1})}$
+$P(x_{t+1},r_{t+1},d|S_{t})$
+$=$ $\sum_{z}\Pr(x_{t+1},r_{t+1},d|z,S_{t})\Pr(z|S_{t})$
+$=$ $\frac{\sum_{z}\Pr(z)\prod_{i=1}^{t+1}\Pr(x_{i},r_{i},d|z,S_{i-1})}{\sum_{z^{\prime}}\Pr(z)\prod_{i=1}^{t}\Pr(x_{i},r_{i},d|z^{\prime},S_{i-1})}.$
 Thus, it holds that
-$-\sum_{t=0}^{T}\log\Pr(x_{t+1},r_{t+1},d|S_{t},K)=$ $-\sum_{t=0}^{T}\begin{aligned} \Big{(}&\log\sum_{z,g}\Pr(z,g|K)\prod_{i=1}^{t+1}\Pr(x_{t},r_{t},d|z,g,S_{t-1},K)\\ &\quad-\log\sum_{z,g}\Pr(z,g|K)\prod_{i=1}^{t}\Pr(x_{t},r_{t},d|z,g,S_{t-1},K)\Big{)}\end{aligned}$
-$=$ $-\log\sum_{z,g}\Pr(z,g|K)\prod_{i=1}^{T+1}\Pr(x_{t},r_{t},d|z,g,S_{t-1},K)$
-$\leq$ $-\log\Pr(z^{*},\dot{g}|K)\prod_{i=1}^{T+1}\Pr(x_{t},r_{t},d|z^{*},\dot{g},S_{t-1},K)$
-$=$ $-\log\Pr(z^{*},\dot{g}|K)-\sum_{i=1}^{T+1}\log\Pr(x_{t},r_{t},d|z^{*},\dot{g},S_{t-1},K)$
-$=$ $-\log\Pr(z^{*},\dot{g}|K)-\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},\dot{g},S_{t-1},K)-\sum_{i=1}^{T}\log\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K).$
+$-\sum_{t=0}^{T}\log\Pr(x_{t+1},r_{t+1},d|S_{t})$
+$=$ $\begin{aligned} &-\sum_{t=0}^{T}\Big(\\ &\log\sum_{z}\Pr(z)\prod_{i=1}^{t+1}\Pr(x_{i},r_{i},d|z,S_{i-1})\\ &-\log\sum_{z}\Pr(z)\prod_{i=1}^{t}\Pr(x_{i},r_{i},d|z,S_{i-1})\\ &\Big)\end{aligned}$
+$=$ $-\log\sum_{z}\Pr(z|K)\prod_{i=1}^{T+1}\Pr(x_{i},r_{t},d|z,S_{i-1})$
+$\leq$ $-\log\Pr(z^{*})\prod_{i=1}^{T+1}\Pr(x_{i},r_{t},d|z^{*},S_{i-1})$
+$=$ $-\log\Pr(z^{*})$
+$-\sum_{i=1}^{T+1}\log\Pr(x_{i},r_{i},d|z^{*},S_{i-1})$
+$=$ $-\log\Pr(z^{*})$
+$-\sum_{i=1}^{T}\log\Pr(r_{i},d|x_{i},z^{*},S_{i-1})$
+$-\sum_{i=1}^{T}\log\Pr(x_{i}|z^{*},S_{i-1}).$
 Thus,
-$-\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1},K)$
-$\leq$ $-\frac{1}{T}\left(\log\Pr(z^{*},\dot{g}|K)+\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},\dot{g},S_{t-1},K)+\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K)}{\Pr(x_{t}|S_{t-1},K)}\right)$
-$+\frac{1}{T}\sum_{i=1}^{T}\log\Pr(d|r_{t},x_{t},S_{t-1},K).$
-$\leq$ $-\frac{1}{T}\left(\log\Pr(z^{*},\dot{g}|K)+\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},\dot{g},S_{t-1},K)+\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K)}{\Pr(x_{t}|S_{t-1},K)}\right)$
-Appendix C Proof of Corollary 3
+$-\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1})$
+$\leq$ $-\frac{1}{T}\bigg(\log\Pr(z^{*})$
+$+\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},S_{t-1})$
+$+\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},S_{t-1})}{\Pr(x_{t}|S_{t-1})})$
+$+\frac{1}{T}\sum_{i=1}^{T}\log\Pr(d|r_{t},x_{t},S_{t-1}\bigg).$
+$\leq$ $-\frac{1}{T}\bigg(\log\Pr(z^{*})$
+$+\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},S_{t-1})$
+$+\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},S_{t-1})}{\Pr(x_{t}|S_{t-1})}\bigg)$
+Appendix C Bounding ICL for Generation
+We can extend the ICL bound shown in Theorem 1 to generation tasks.
+When using ICL for generation, in the demonstration part of the prompt, people usually include a separator between the input and the output.
+For example, if the task is to translate a sentence to English, people may write the demonstration in the format “[input] should be converted to [output].” This “should be converted to” is an expression that can be associated to the “meaning” of the task, namely “should be mapped to a translation in English.” The intuition of the theorem is that, a language model may be able to uncover this association through the demonstration examples via the same mechanism as how it perform ICL for classification tasks.
+Theorem C.1 (Average ICL Loss for Generation).
+Let $z^{*}$ represent the association between a separator $\xi$ and a task descriptions such that for any input $x$, $P(x,\xi,r|z^{*})>0$ only if $r$ is one of the correct outputs as specified by the task.
+Let $K$ be the constraints used for decoding, and $\dot{g}$ be the event where a document follows certain formats.
+Let $R_{t}$ be the set of correct outputs for $x_{t}$, $S_{t}=\{x_{1},\xi,r_{2},d,x_{2},\xi,r_{2},\cdots,x_{t},\xi,r_{t},d|r_{1}\in R_{1},r_{2}\in R_{2},\cdots,r_{t}\in R_{t}\}$.
+We have for any integer $T>0$, the average cross-entropy loss of ICL is bounded as:
+$\begin{split}&-\frac{1}{T}\sum_{t=0}^{T}\log\Pr(R_{t}|x_{t},\xi,S_{t-1})\\ &\leq\begin{aligned} &-\frac{1}{T}\log\Pr(z^{*})\\ &-\frac{1}{T}\sum_{i=1}^{T}\log\Pr(R_{t},d|x_{t},\xi,z^{*},S_{t-1})\\ &-\frac{1}{T}\sum_{i=1}^{T}\log\frac{\Pr(x_{t},\xi|z^{*},S_{t-1})}{\Pr(x_{t},\xi|S_{t-1})}\end{aligned}\end{split}$ (4)
+Proof.
+If we see $\xi$ as a part of the $x$’s in the proof for Theorem 1, then following the steps in §B, we can have Eq. 4. ∎
+Appendix D Proof of Corollary 2
 The second term in the right-hand side of Eq. 1 is zero when the decoding constrain $K$ is imposed. Therefore, it suffices to prove the last term is non-negative in expectation.
-$\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K)}{\Pr(x_{t}|S_{t-1},K)}=$ $\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},K)}{\Pr(x_{t}|S_{t-1},K)}$
-$=$ $\sum_{x_{1},x_{2},\cdots,x_{T}}\Pr(x_{t}|z^{*},\dot{g},K)\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},K)}{\Pr(x_{t}|S_{t-1},K)}$
-$=$ $\mathrm{KLD}(\Pr(x_{t}|z^{*},\dot{g},K)||\Pr(x_{t}|S_{t-1},K))\geq 0$
-Appendix D The Connection between $P(r_{t}|x_{t},z^{*},\ddot{g})$ and Function Description Length by Hahn & Goyal (2023)
+$\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},S_{t-1},K)}{\Pr(x_{t}|S_{t-1},K)}$
+$=$ $\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},K)}{\Pr(x_{t}|S_{t-1},K)}$
+$=$ $\sum_{x_{1},x_{2},\cdots,x_{T}}\Pr(x_{t}|z^{*},K)\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},K)}{\Pr(x_{t}|S_{t-1},K)}$
+$=$ $\mathrm{KLD}(\Pr(x_{t}|z^{*},K)||\Pr(x_{t}|S_{t-1},K))\geq 0$
+Appendix E The Connection between $P(r_{t}|x_{t},z^{*})$ and Function Description Length by Hahn and Goyal (2023)
 Firstly, we make some regularity assumptions: Given a step-by-step reasoning process $\pi=s_{1},s_{2},\cdots,s_{n}$ for the induction process of $P\models Q$, in the training data,
@@
 We first show how we derive Eq. 3: Based on Assumption 6.1,
-$\Pr(r_{t}|x_{t},z^{*},\ddot{g})=\sum_{\pi\in\Pi}\Pr(\pi,r_{t}|x_{t},z^{*},\ddot{g})\Pr(\text{$\pi$ is dropped}),$
+$\begin{split}&\Pr(r_{t}|x_{t},z^{*})\\ &=\sum_{\pi\in\Pi}\Pr(\pi,r_{t}|x_{t},z^{*})\Pr(\text{$\pi$ is dropped}),\end{split}$
 where $\Pi$ is a set of token sequences representing reasoning steps that induce $r_{t}$ from $x_{t}$. Let $\pi^{*}$ be the shortest proof in $\Pi$, we have
-$\log\Pr(r_{t}|x_{t},z^{*},\ddot{g})=$ $\log\sum_{\pi\in\Pi}\Pr(\pi,r_{t}|x_{t},z^{*},\ddot{g})\Pr(\text{$\pi$ is dropped})$
-$\geq$ $\log\Pr(\pi^{*},r_{t}|x_{t},z^{*},\ddot{g})\Pr(\text{$\pi^{*}$ isdropped})$
+$\log\Pr(r_{t}|x_{t},z^{*})$
+$=$ $\log\sum_{\pi\in\Pi}\Pr(\pi,r_{t}|x_{t},z^{*})\Pr(\text{$\pi$ is dropped})$
+$\geq$ $\log\Pr(\pi^{*},r_{t}|x_{t},z^{*})\Pr(\text{$\pi^{*}$ is dropped})$
 $\geq$ $p_{min}\log\ell(\pi^{*})+p_{drop}\log\ell(\pi^{*}).$
-Then we can discuss the connection between $\Pr(r_{t}|x_{t},z^{*},\ddot{g})$ and the function description length by Hahn & Goyal (2023).
-We can see the dropped reasoning steps in $\pi^{*}$ as the hidden (tree) structure that maps $x_{t}$ to $r_{t}$ as the derivation tree $\tau_{\phi}$ in the bound of Hahn & Goyal (2023).
+Then we can discuss the connection between $\Pr(r_{t}|x_{t},z^{*},\ddot{g})$ and the function description length by Hahn and Goyal (2023).
+We can see the dropped reasoning steps in $\pi^{*}$ as the hidden (tree) structure that maps $x_{t}$ to $r_{t}$ as the derivation tree $\tau_{\phi}$ in the bound of Hahn and Goyal (2023).
 The length of the reasoning steps thus corresponds to the description length of the derivation tree $\mathrm{D}(\tau_{\phi})$.
-A major difference between the bound by Hahn & Goyal (2023) and our bound is that their bound has $\mathrm{D}(\tau_{\phi})$ constant to $T$ while our bound has $\sum_{t}\log\Pr(r_{t}|x_{t},z^{*},\ddot{g})$, which potentially grows proportionally to $T$.
-The cause of this difference is that, Hahn & Goyal (2023) assumes a structure that repetitively applies a function mapping in a document, and the number of repetition is independent to the complexity of the function mapping.
+A major difference between the bound by Hahn and Goyal (2023) and our bound is that their bound has $\mathrm{D}(\tau_{\phi})$ constant to $T$ while our bound has $\sum_{t}\log\Pr(r_{t}|x_{t},z^{*},\ddot{g})$, which potentially grows proportionally to $T$.
+The cause of this difference is that, Hahn and Goyal (2023) assumes a structure that repetitively applies a function mapping in a document, and the number of repetition is independent to the complexity of the function mapping.
 In comparison, our framework does not make this assumption.
-Appendix E Detailed Gengeration Process of the LM Training Data in Calcutec
+Appendix F Details of the Calcutec Experiment
+F.1 Generation Process of the LM Training Data in Calcutec
 We generate a paragraph based on Assumption 3.3 in the following step:
@@
    if $|M^{\prime}|>\frac{|\Sigma|}{2}$ then
-           Reject the sampled $X\cup\{x^{\prime}\}$. {We don’t want a premise that entails everything.}
+     Reject the sampled $X\cup\{x^{\prime}\}$. $\triangleright$ We don’t want a premise that entails everything.
      Restart the while loop.
    end if
-        Sample a $g\in M^{\prime}$ such that the proof tree for $X^{\prime}\models g$ contains $s$ and its depth $>d_{min}$. {We use $d_{min}=4$ in our experiments.}
+   Sample a $g\in M^{\prime}$ such that the proof tree for $X^{\prime}\models g$ contains $s$ and its depth $d_{min}$. $\triangleright$ We use $d_{min}=4$ in our experiments.
    Do topological sort to flatten the proof tree and convert it into a string.
@@
 for $s\in S$ do
-     $D\leftarrow$ $D$.replace($s$, $r_{a}$)
+  $\leftarrow DD$.replace($s$, $r_{a}$)
 end for
@@
 for $s^{\prime}\in S^{\prime}$ do
-     $D\leftarrow$ $D$.replace($s^{\prime}$, $r_{b}$)
+  $\leftarrow DD$.replace($s^{\prime}$, $r_{b}$)
 end for
-Algorithm 1 Pseudo code for the generation process of an Calcutec document used for training.
-Appendix F Perturbations in Calcutec
-We apply two types of perturbations over the reasoning steps in Calcutec described in §4:
+Algorithm 1 Pseudo code for the generation process of a Calcutec document used for training.
+F.2 Perturbations in Calcutec
+We apply two types of perturbations over the reasoning steps in Calcutec described in §7.1:
   1. 1.
@@
 We use $p_{merge}=p_{drop}=p_{skip}$.
-Additionally, when flatting the proof trees with topological sort, we break ties randomly. We also randomize the order of the symbols in the antecedents.
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Additionally, when flattening the proof trees with topological sort, we break the ties randomly. We also randomize the order of the symbols in the antecedents.
+(a) document (tokens)
+(b) paragraph (tokens)
+(c) paragraph (steps)
+(d) first $z$ (steps)
+(e) document (tokens)
+(f) paragraph (tokens)
+(g) paragraph (steps)
+(h) first $z$ (steps)
 Figure 3: The distribution of lengths and the first step in each paragraph where $z$ is the consequence in the Calcutec dataset.
 The first/second row are the statistics before/after some steps are randomly dropped.
-Appendix G Extra Experiments with Calcutec
-G.1 Additional Setups
+F.3 Extra Setups
 Unseen Inference Process.
-Based on Assumption 2.9 and Assumption 2.5, input-label pairs of a downstream task corresponds to prefix-reference pairs in a paragraph.
+Based on Assumption 6.1 and the formalism of NLP tasks in §3.2, input-label pairs of a downstream task corresponds to prefix-reference pairs in a paragraph.
 To examine whether the trained model can generalize well when the induction process for the label is different from the induction process for the pronoun in the training data, we generate a training dataset where all the pronouns are induced from the premise with a left-branching proof tree with a depth equal to 2 (Figure 4(a)), while the test data contains samples whose labels are induced from the input with balanced trees (Figure 4(b)).
@@
 No perturbations.
-As described in §F, we apply some random perturbations on the proving process. We also experiment with the setup where we do not apply any perturbations.
+As described in §F.2, we apply some random perturbations on the proving process. We also experiment with the setup where we do not apply any perturbations.
 With/Without Rewriting the First Step.
-As described in §E, we rewrite the first step of the proof. We also experiment with the setup where we do not rewrite the first step.
+As described in §F.1, we rewrite the first step of the proof. We also experiment with the setup where we do not rewrite the first step.
 Model Size.
 We also experiment with different models sizes. We experiment with GPT-2 models that have 3, 4 and 5 layers.
-G.2 Results and Discussion
+F.3.1 Results and Discussion
 Unseen Inference Process.
@@
 Model Size.
-Figure 8 show that deeper models have better ICL performance. It is aligned with the real-world observation that scaling helps model performance.
-[…]
-[…]
+Figure 9 show that deeper models have better ICL performance. It aligns with the real-world observation that scaling helps model performance.
+                                      $x_{1}$$x_{2}$$x_{4}$ $x_{3}$$r_{1}$ $x_{1}$$x_{3}$$x_{6}$ $x_{7}$ $\vdots\vdots$
 (a) The proof tree a paragraph in the training dataset corresponds .
-[…]
-[…]
+                       $x_{1}$$x_{2}$$x_{5}$ $x_{3}$$x_{4}$$x_{6}$ $r_{1}$
 (b) A balanced tree for a downstream task sample.
 Figure 4: Proof trees examples.
-Appendix H Hyper-parameters
+F.4 Hyper-parameters
 We train our model using batch size 256, warm up ratio 5%, and we truncate the sequence length to 512 tokens and the default parameters for the optimizer.
@@
 All models can be trained with 4 RTX 2080ti within 8 hours.
-(a) balanced, length = $4$, $y_{1},y_{2}$
-(b) branching, length = $4$, $y_{1},y_{2}$
-(c) branching, length = $3$, $y_{1},y_{2}$
-(d) balanced, length = $3$, $y_{3},y_{4}$
-(e) branching, length = $4$, $y_{3},y_{4}$
-(f) branching, length = $3$, $y_{3},y_{4}$
+(a) balanced, length = 44, $y_{1},y_{2}$
+(b) branching, length = 44, $y_{1},y_{2}$
+(c) branching, length = 33, $y_{1},y_{2}$
+(d) balanced, length = 33, $y_{3},y_{4}$
+(e) branching, length = 44, $y_{3},y_{4}$
+(f) branching, length = 33, $y_{3},y_{4}$
 Figure 5: In-context learning accuracy with Calcutec when using different verbalizers ($y_{1},y_{2}$ or $y_{3},y_{4}$) and input lengths (3 or 4).
-The dotted lines represent the performance of unseen combinations described in §4.2, while the different colors represent the number of formulas each class ($v_{+}$ or $v_{-}$) is associated to.
+The dotted lines represent the performance of unseen combinations described in §7.1.2, while the different colors represent the number of formulas each class ($v_{+}$ or $v_{-}$) is associated to.
 The main lines represent the average accuracy of 5 tasks.
 We plot the performance of each task in lighter colors.
-(a) balanced, length = $4$, $r_{1},r_{2}$
-(b) branching, length = $4$, $r_{1},r_{2}$
-(c) branching, length = $3$, $r_{1},r_{2}$
-(d) balanced, length = $4$, $r_{3},r_{4}$
-(e) branching, length = $4$, $r_{3},r_{4}$
-(f) branching, length = $3$, $r_{3},r_{4}$
+(a) balanced, length = 44, $r_{1},r_{2}$
+(b) branching, length = 44, $r_{1},r_{2}$
+(c) branching, length = 33, $r_{1},r_{2}$
+(d) balanced, length = 44, $r_{3},r_{4}$
+(e) branching, length = 44, $r_{3},r_{4}$
+(f) branching, length = 33, $r_{3},r_{4}$
 Figure 6: In-context learning accuracy with Calcutec when no steps are dropped ($p_{skip}=0$).
-(a) balanced, length = $4$, $r_{1},r_{2}$
-(b) branching, length = $4$, $r_{1},r_{2}$
-(c) branching, length = $3$, $r_{1},r_{2}$
-(d) balanced, length = $4$, $r_{3},r_{4}$
-(e) branching, length = $4$, $r_{3},r_{4}$
-(f) branching, length = $3$, $r_{3},r_{4}$
+(a) balanced, length = 44, $r_{1},r_{2}$
+(b) branching, length = 44, $r_{1},r_{2}$
+(c) branching, length = 33, $r_{1},r_{2}$
+(d) balanced, length = 44, $r_{3},r_{4}$
+(e) branching, length = 44, $r_{3},r_{4}$
+(f) branching, length = 33, $r_{3},r_{4}$
 Figure 7: In-context learning accuracy with Calcutec without rewriting the first step to include contain the premise of the proof.
-(a) balanced, length = $4$, 3 layers
-(b) balanced, length = $4$, 4 layers
-(c) balanced, length = $4$, 5 layers
-(d) branching, length = $4$, 3 layers
-(e) branching, length = $4$, 4 layers
-(f) branching, length = $4$, 5 layers
+(a) balanced, $r_{1},r_{2}$, 25000 examples
+(b) balanced, $r_{1},r_{2}$, 10000 examples
+(c) balanced, $r_{1},r_{2}$, 5000 examples
+(d) balanced, $r_{3},r_{4}$, 25000 examples
+(e) balanced, $r_{3},r_{4}$, 10000 examples
+(f) balanced, $r_{3},r_{4}$, 5000 examples
+(g) branching, $r_{1},r_{2}$, 25000 examples
+(h) branching, $r_{1},r_{2}$, 10000 examples
+(i) branching, $r_{1},r_{2}$, 5000 examples
+(j) branching, $r_{3},r_{4}$, 25000 examples
+(k) branching, $r_{3},r_{4}$, 10000 examples
+(l) branching, $r_{3},r_{4}$, 5000 examples
+Figure 8: In-context learning accuracy with different sizes of
+(a) balanced, length = 44, 3 layers
+(b) balanced, length = 44, 4 layers
+(c) balanced, length = 44, 5 layers
+(d) branching, length = 44, 3 layers
+(e) branching, length = 44, 4 layers
+(f) branching, length = 44, 5 layers
 Figure 9: The in-context learning performance when using models with different model depths.
@@
 Table 5: The 4-shot accuracy of in-context learning (ICL) versus chain-of-thoughts (CoT).
-(a) 5-digit
-(b) 6-digit
-Figure 9: The distribution of the number of reasoning steps in the dataset when some of them are dropped at different probability.
-Each number is the average over 5 datasets generated with different random seeds.
-Appendix I Formal Description of the Digit Addition Data
-[…]
-(a) 5 digits
-(b) 6 digits
-Figure 10: The accuracy of the models for the addition tasks.
-The x-axis represents the probability at which we drop each reasoning step in the training data independently.
-The solid line represents the ratio of testing samples where the model can output the exact answer, while the dashed line represents the character-level accuracy.
-The results are the average of 5 random seeds.
-Figure 11: The exact accuracy (y-axis, solid points) and digit-level accuracy (y-axis, hollow points) versus validation loss (x-axis) for a 5-digit addition task.
-Figure 12: The exact accuracy (y-axis, solid points) and digit-level accuracy (y-axis, hollow points) versus validation loss (x-axis) for a 6-digit addition task.
+branching balance
+$r_{1},r_{2}$ $r_{3},r_{4}$ $r_{1},r_{2}$ $r_{3},r_{4}$
+#-shot 2 4 6 2 4 6 2 4 6 2 4 6
+single 49.1 89.5 84.0 59.5 92.0 86.9 58.5 86.2 85.5 50.3 90.3 89.9
+double 47.8 71.4 75.6 53.1 77.1 86.1 49.1 70.4 69.0 50.5 75.8 79.4
+triple 46.7 65.7 70.7 50.6 73.4 79.4 46.0 60.2 61.4 49.8 67.0 70.4
+Table 6: The CoT performance with 2, 4, or 6 examples.
+Sample $N=\{n_{1},n_{2},n_{3},n_{4},n_{5}\}$ from either $\{0,1,\cdots,9\}$ or $\{6,7,\cdots,15\}$.
+Sample four operators $O=\{o_{1},o_{2},o_{3},o_{4}\}$ from $\{+,\times\}$.
+Initialize an empty list $S$ for the storing the steps.
+while $\|N\|>1$ do
+  Randomly sample an index $i$ of an operator from $O$. If $\times\in O$, then make sure that the sampled index corresponds to $\times$.
+  Create a step $s\leftarrow[n_{i},n_{i+1},n_{i}o_{i}n_{i+1}\mod 16]$ and append $s$ to $S$.
+  Remove $n_{i}$ and $n_{i+1}$ from $N$.
+  Remove $o_{i}$ from $O$.
+  Insert $n_{i}o_{i}n_{i+1}\mod 16$ into NN at position ii.
+end while
+return $S$.
+Algorithm 2 Pseudo code for the generation process of a training example for the arithmetic task.
+Given a list of steps $S=\langle s_{1},s_{2},\cdots,s_{4}\rangle$. (Each step is represented with a list, where the last number in the list is the right-hand side of an equation.)
+Initialize n list $S^{\prime}=\langle s_{1}\rangle$ for the storing the randomly merged steps.
+for $i\in\{2,3,4,5\}$ do
+  Uniformly sample a number $r$ from $\{0,1\}$.
+  if $r=0$ then
+   Append $s_{i}$ to $S^{\prime}$.
+  else
+   if $S^{\prime}[-1][-1]\in s_{i}[:-1]$ then
+     Remove $S^{\prime}[-1][-1]$ from $s_{i}$.
+     $S^{\prime}[-1]\leftarrow\text{concatenate}(S^{\prime}[-1][:-1],s_{i})$
+   end if
+  end if
+end for
+return $S^{\prime}$.
+Algorithm 3 Pseudo code for the generation process of a training example for the arithmetic task.
+Appendix G Details of the Arithmetic Task
+G.1 Training Data
+When generating a training example, we first sample five numbers.
+These five numbers are either from $\{0,\cdots,9\}$ or $\{6,\cdots,15\}$.
+We then repeat the following process until there is only one number remaining in the list: We randomly pick two neighbor numbers from the list of five numbers and sum them and insert the result back into the list.
+This process results in a sequence of four steps, each of which is an equation that adds two numbers and produces one number.
+We outline the algorithm in Algorithm 2.
+After we generate the four steps, we randomly merge steps into one step.
+We iterate through the steps.
+If the left-hand side current step contains the result of a previous step, we merge these two step into a single step at probability 0.5.
+For example, if the two steps are “1 + 2 = 3 .
+3 + 4 = 7.” We merge these two steps as “1 + 2 + 4 = 7.” We outline the algorithm in Algorithm 3.
+G.2 Hyper-parameters
+We generate 80000 examples for training, 10000 for validation and 10000 for evaluation in the seen setup.
+For the unseen setup, we generate $6^{5}=7776$ examples.
+We train five Transformer models of the GPT-2-small architecture with five random seeds for 15 epochs.
+We use batch size 64 and the default training hyper parameters in HuggingFace transformers-4.46.3, i.e., with learning rate 5e-5, warm-up ratio 0, AdamW optimizer.
+Appendix H Dataset License
+  • SST-2: MIT
+  • MR: unavailable
+  • CR: unavailable
+  • Subj: unavailable